### v1

# Finite-Time Analysis of Gradient Descent for Shallow Transformers

Abstract
Understanding why Transformers perform so well remains challenging due to their non-convex optimization landscape. In this work, we analyze a shallow Transformer with $m$ independent heads trained by projected gradient descent in the kernel regime. Our analysis reveals two main findings: (i) the width required for nonasymptotic guarantees scales only logarithmically with the sample size $n$, and (ii) the optimization error is independent of the sequence length $T$. This contrasts sharply with recurrent architectures, where the optimization error can grow exponentially with $T$. The trade-off is memory: to keep the full context, the Transformer’s memory requirement grows with the sequence length. We validate our theoretical results numerically in a teacher–student setting and confirm the predicted scaling laws for Transformers.
1 INTRODUCTION
Transformers have reshaped modern machine learning, achieving state-of-the-art results in language, vision, and multi-modal tasks (Vaswani et al., 2017; Devlin et al., 2019; Brown et al., 2020; Dosovitskiy et al., 2021). Despite this empirical success, a sharp nonasymptotic understanding of their training dynamics remains elusive. Two obstacles stand out: (i) the highly nonlinear nature of the self-attention layer, whereby each output coordinate depends on all inputs through a softmax reweighting, making standard arguments for fully-connected networks or CNNs difficult to apply; and (ii) the nonconvexity of the training landscape, which complicates guarantees for gradient-based methods outside of infinite-width limits.
A growing theory literature has therefore analyzed simplified settings. To make comparisons precise, let us recall the single-head self-attention layer. For a length-$T$ input $X\in\mathbb{R}^{d\times T}$ and weights $W_{Q},W_{K},W_{V}\in\mathbb{R}^{m\times d}$, the attention layer outputs
$a(X)\;=\;W_{V}X\,\sigma_{s}\!\big(\beta X^{\top}W_{K}^{\top}W_{Q}X\big),$ (1)
where $\sigma_{s}$ denotes the column-wise softmax, and $\beta$ is the softmax scaling. In the original Transformer, one takes $\beta=m^{-1/2}$ to stabilize the dot-product variance (Vaswani et al., 2017). However, to make the analysis of (1) tractable, prior work often studies variants that take one of the following three directions: (i) avoid $\sigma_{s}$ entirely to obtain linear attention (Katharopoulos et al., 2020; Choromanski et al., 2021; Lu et al., 2021; Zheng et al., 2022; Zhang et al., 2024); (ii) replace $\sigma_{s}$ with a pointwise activation such as ReLU (Wortsman et al., 2023; Hron et al., 2020); or (iii) take infinite-width limits under $\beta=m^{-1}$, in which the softmax weights approach uniform distribution and the attention layer reduces to mean-pooling (Wu et al., 2023; Noci et al., 2023; Bordelon et al., 2024). While analytically convenient, these modifications limit the scope of the resulting guarantees.
More recently, an interesting work by Wu et al. (2023) established global convergence for shallow Transformers at finite width under the realistic $\beta=m^{-1/2}$ scaling, showing that cubic overparameterization in the sample size $n$ suffices for gradient descent to converge at a linear rate. Their proof follows the usual neural tangent kernel (NTK) recipe of enforcing a strict positive definiteness to a Gram matrix to drive convergence (Du et al., 2019a, b; Allen-Zhu et al., 2019; Zou and Gu, 2019; Nguyen, 2021; Nguyen et al., 2021). In particular, Wu et al. (2023) control the minimum singular value of a feature matrix built from the last hidden layer, which relies on the regime $d\geq n$, thereby limiting its practicality.
Parallel to these developments for Transformers, new techniques have sharpened finite-time analyses for recurrent architectures (Cayci and Eryilmaz, 2024a, b). Cayci and Eryilmaz (2024a) prove nonasymptotic convergence of gradient descent for IndRNNs with dramatically improved width requirements—most notably, width scaling only logarithmically in the sample size $n$. Their proof is organized around norm-constrained transportation mappings that represent the target dynamical system in the RKHS induced by the RNN’s NTK, tying random-feature ideas to near-initialization training (Rahimi and Recht, 2007; Ji et al., 2020). Together with a projection step that keeps the gradient-descent iterates near initialization, they obtain finite-sample, finite-width convergence. The bounds also make explicit the price of long-term dependencies: beyond a cutoff governed by the activation’s Lipschitz constant, both iteration complexity and the required width can grow exponentially with the sequence length $T$.
In this paper, building on techniques from Cayci and Eryilmaz (2024a), we analyze a shallow multi-head Transformer trained by projected gradient descent algorithms (ProjGD and ProjSGD) in the near-initialization regime while preserving the nonlinearity of the attention layer. We address the nonlinearity and nonconvexity of the problem by working with projected updates that keep parameters in a small neighborhood of initialization. In this regime, a transportation-mapping viewpoint links the random features model induced at initialization to the NTK linearization and provides uniform control of the linearization error. We then exploit an explicit gradient characterization of the attention pathway to obtain optimization bounds independent of sequence length $T$. The main contributions of this work can be highlighted as follows:
•  Preserving attention nonlinearity with multiple heads. Unlike the approaches discussed above that remove or linearize softmax, or that enforce degeneracy via the scaling $\beta=m^{-1}$, our analysis retains the genuine attention nonlinearity and allows independent heads.
•  Bypassing positive definiteness of the NTK. Unlike the original works in the NTK regime, our analysis does not require strict positive definiteness of the NTK.
•  Logarithmic overparameterization. We show that, in our setting, the width required for nonasymptotic training guarantees scales only logarithmically with the sample size $n$. To the best of our knowledge, this has not been shown for Transformers.
•  Optimization error independent of input sequence length $T$. Our ProjGD bound does not degrade with the sequence length $T$. This contrasts sharply with RNN analyses, where the dependence on $T$ can be exponential under long-term dependencies. The trade-off is memory: retaining full context implies that the Transformer’s memory footprint grows with $T$ due to attention over all tokens.
•  Experimental validation. With controlled experiments in teacher–student settings, we validate the predicted scaling laws for our Transformer model and contrast Transformers with IndRNNs in their handling of long-term dependencies and memory complexity.
2 RELATED WORK
Neural Tangent Kernel (NTK) analysis.
The NTK framework linearizes networks around initialization and connects gradient descent to kernel gradient flow, yielding optimization and generalization guarantees in overparameterized regimes (Jacot et al., 2018; Lee et al., 2019; Du et al., 2019b; Chizat et al., 2019). For CNNs, NTK limits capture locality and weight sharing (Garriga-Alonso et al., 2019; Novak et al., 2019), and the convolutional NTK admits exact computation (Arora et al., 2019). For RNNs, the Recurrent NTK defines sequence kernels that aggregate contributions across time (Alemohammad et al., 2021).
Asymptotic theory for Transformers.
Representation‑theoretic results establish that Transformers are universal approximators of permutation‑equivariant sequence‑to‑sequence maps, and, with trainable positional encodings, of general continuous sequence‑to‑sequence functions on compact domains (Yun et al., 2020). At infinite width, attention admits NTK descriptions under suitable scalings: Hron et al. (2020) derive attention NTK limits, showing that under the common $\beta=m^{-1/2}$, letting the number of heads grow results in a deterministic NTK. Under $\beta=m^{-1}$, deterministic kernels can exist but may degenerate toward mean-pooling or require $W_{Q}=W_{K}$ (Yang, 2020). Wu et al. (2023) complement their finite‑width convergence proof with an NTK analysis for shallow Transformers with $\beta=m^{-1}$.
Nonasymptotic theory for Transformers.
Wu et al. (2023) prove global convergence of a shallow Transformer while preserving the softmax attention nonlinearity and carefully diagnosing scaling/initialization choices. Complementary results capture in‑context learning dynamics with true softmax: Huang et al. (2024) give finite‑time guarantees and a phase‑wise view of attention; Song et al. (2024) show convergence under suitable initialization and reveal that softmax can introduce spurious local minima whereas a Gaussian attention smooths the landscape. Simplified settings using linear attention establish convergence for linear predictors learned in‑context (Zhang et al., 2024). Beyond linear tasks, Shen et al. (2024) obtain linear‑rate convergence to the Bayes classifier for in‑context classification with a shallow multi‑head Transformer. Further optimization views sharpen what attention learns: Tarzanagh et al. (2023) show that gradient descent on a single attention layer aligns with a max‑margin (SVM‑like) token‑selection rule. Nichani et al. (2024) analyze a two‑layer attention‑only Transformer and prove that gradient descent recovers the latent causal structure of the in-context learning problem.
Notation.
For $k\in\mathbb{N}$ we write $k]=\{1,2,\dots,k\}$. For matrices $A,B$ we write $\langle A,B\rangle_{F}=\mathrm{tr}(A^{\top}B)$ and $\|A\|_{F}$ for the Frobenius norm. The probability simplex in $\mathbb{R}^{T}$ is $\Delta_{T}=\{\,\alpha\in\mathbb{R}^{T}:\alpha_{t}\geq 0,\ \sum_{t=1}^{T}\alpha_{t}=1\,\}$ and $\mathbb{R}_{+}=(0,\infty)$ is the set of positive reals. For a vector $x$, $\operatorname{diag}(x)$ denotes the diagonal matrix with diagonal $x$. For matrices $A_{1},\ldots,A_{r}$, $\operatorname{blkdiag}(A_{1},\ldots,A_{r})$ denotes the block-diagonal matrix with diagonal blocks $A_{1},\ldots,A_{r}$ and zero elsewhere. For $A\in\mathbb{R}^{m\times n}$, $\operatorname{vec}(A)\in\mathbb{R}^{mn}$ denotes the column-wise vectorization of $A$. The notation $\lesssim_{a,b}$ omits the constants depending on the parameters $a,b$. For vectors $x,y$, $x\leq y$ denotes element-wise comparison. $\mathrm{Unif}(V)$ denotes the uniform distribution over the set $V$, and $\mathrm{Rad}(\alpha)$ denotes $\mathrm{Unif}(\{-\alpha,\alpha\})$.
3 PROBLEM SETUP
We specify the objective, introduce the shallow multi-head Transformer studied in this work, and formalize the supervised learning task with the optimization algorithms that will be analyzed.
3.1 Data
Each sample $X=(X_{1},\dots,X_{T})\in\mathbb{R}^{d\times T}$ is a sequence of $d$-dimensional tokens coming from a compact set
$\mathcal{X}\;\coloneqq\;\bigl\{\,X\in\mathbb{R}^{d\times T}\;:\;\max_{t\in[T]}\|X_{t}\|_{2}\leq 1\,\bigr\}.$
The label is a scalar $y\in\mathbb{R}$. We observe samples from a dataset
$\mathcal{D}=\bigl\{(X^{(j)},y^{(j)})\bigr\}_{j=1}^{n}\ \subset\ \mathcal{X}\times\mathbb{R}$
that were generated by a function $f^{\star}$. In other words, $f^{\star}(X^{(j)})=y^{(j)}$ for every $j\in[n]$. Our goal will be to approximate the function $f^{\star}$ with the Transformer model defined in the next section.
3.2 Transformer Architecture
Fix $X\in\mathcal{X}$. Since each training sample has a scalar target, we employ pooled self-attention driven by a single query $q_{X}\in\mathbb{R}^{d}$ with $\|q_{X}\|_{2}\leq 1$. The query $q_{X}$ is not trainable and may be a fixed vector ([CLS] as in BERT (Devlin et al., 2019), or ViT’s class token (Dosovitskiy et al., 2021)), a deterministic function of the input (e.g., last-token pooling $q_{X}=X_{T}$ for next-token prediction (Radford et al., 2019)), or a pre-specified seed query used to summarize high-dimensional inputs (Jaegle et al., 2021).
For $i\in[m]$ and trainable $W_{i}\!\in\!\mathbb{R}^{d\times d}$, $U_{i}\!\in\!\mathbb{R}^{d}$, and $c_{i}\!\in\!\mathbb{R}$, we consider the following model.
•  Attention layer.
$a(X;W_{i})\coloneqq X\,\sigma_{s}\!\big(X^{\top}W_{i}q_{X}\big)\in\mathbb{R}^{d},$ (2)
where $\sigma_{s}:\mathbb{R}^{T}\!\to\!\Delta_{T}$ is the softmax:
$(\sigma_{s}(z))_{t}\;=\;\frac{\exp(z_{t})}{\sum_{s=1}^{T}\exp(z_{s})}.$
•  Feed-forward layer.
$h(X;\theta_{i})\coloneqq\sigma\!\big(U_{i}^{\top}a(X;W_{i})\big)\in\mathbb{R},$ (3)
where $\theta_{i}\coloneqq(U_{i},\operatorname{vec}(W_{i}))$ and the activation $\sigma:\mathbb{R}\!\to\!\mathbb{R}$ satisfies uniform bounds
$|\sigma(u)|\leq\sigma_{0},\;|\sigma^{\prime}(u)|\leq\sigma_{1},\;|\sigma^{\prime\prime}(u)|\leq\sigma_{2}\quad\forall u\in\mathbb{R}.$
We collect these bounds as $\bar{\sigma}=(\sigma_{0},\sigma_{1},\sigma_{2})\in\mathbb{R}^{3}_{+}$.
•  Linear layer.
$f(X;\varphi)\coloneqq\tfrac{1}{\sqrt{m}}\sum_{i=1}^{m}c_{i}\,h(X;\theta_{i})\in\mathbb{R},$ (4)
where $\varphi_{i}\coloneqq(c_{i},U_{i},\operatorname{vec}(W_{i}))$ and $\varphi:=(\varphi_{1},\dotsc,\varphi_{m})$.
Remark 1.
Following prior analyses (Wu et al., 2023; Tarzanagh et al., 2023; Nichani et al., 2024; Oymak et al., 2023) we combine the query and key matrices in (1) into a single attention weight matrix $W_{i}$, which corresponds to the canonical $\beta=m^{-1/2}$ scaling, and absorb the value matrix into the next hidden layer’s weights $U_{i}$.
Remark 2.
In standard multi-head attention one concatenates $h$ attention heads into $A(X)=(a(X;W_{1}),\ldots,a(X;W_{h}))\in\mathbb{R}^{hd}$ and applies a feed-forward layer with a weight matrix $V\in\mathbb{R}^{m\times hd}$, resulting in $\sigma(VA(X))$. Our architecture is equivalent to this with a block-diagonal weight matrix with one neuron per head, i.e., $h=m$ and $V=\operatorname{blkdiag}(U_{1}^{\top},\ldots,U_{m}^{\top})$. This independent head constraint acts as a structural regularizer: it preserves per-head interpretability, reduces cross-head interference, restores a deterministic NTK, and simplifies the analysis without requiring any of the architectural changes mentioned in Section 1.
Remark 3.
A key point to establish our main results will be to show the local Lipschitz continuity and smoothness of the mapping $\theta_{i}\mapsto h(X;\theta_{i})$ in Lemma 2.
3.3 Algorithms for Empirical Risk Minimization
For a given dataset $\mathcal{D}$, parameters $\varphi$ and the predictor $f(\cdot;\varphi)$, we analyze the empirical mean-squared error (MSE) loss, which is expressed as
$\widehat{\mathcal{L}}_{n}(\varphi):=\frac{1}{n}\sum_{j=1}^{n}\bigl(f(X^{(j)};\varphi)-y^{(j)}\bigr)^{2}.$
We optimize $\widehat{\mathcal{L}}_{n}$ by projected methods onto the feasible set $\Omega$. Let $\ell_{j}(\varphi):=\big(f(X^{(j)};\varphi)-y^{(j)}\big)^{2}$ and, for any index set $B\subseteq[n]$, $\widehat{\mathcal{L}}_{B}(\varphi)\;:=\;\frac{1}{|B|}\sum_{j\in B}\ell_{j}(\varphi).$ Given a step size $\eta$ and mini‑batches $\{B_{s}\}_{s\geq 0}$, the iterate updates are, for $s=0,1,2,\dots$:
$\varphi^{(s+1)}$ $:=\Pi_{\Omega}\left(\varphi^{(s)}-\eta\,\nabla_{\varphi}\widehat{\mathcal{L}}_{B_{s}}\!\big(\varphi^{(s)}\big)\right),$
We analyze (i) ProjGD with $B_{s}=[n]$ for all $s$, and (ii) ProjSGD with $B_{s}=\{J_{s}\}$ with $J_{s}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathrm{Unif}([n])$.
4 OVERVIEW OF MAIN RESULTS
Below we present our main results informally, which are formally established with explicit constants in Section 5.5.
(Informal) Theorem 1.
Assuming the target function $f^{\star}$ belongs to the sup-norm constrained Hilbert Space $\mathcal{F}_{\bar{\nu}}$ (formalized in Section 5.4) and we run ProjGD for $\tau$ steps, the average iterate $\bar{\varphi}^{(\tau)}:=\frac{1}{\tau}\sum_{s=0}^{\tau-1}\varphi^{(s)}$ satisfies
$\widehat{\mathcal{L}}_{n}\big(\bar{\varphi}$ $^{(\tau)}\big)\lesssim\underbrace{\tfrac{(d+\log(m/\delta^{\prime}))^{2}}{\sqrt{\tau}}}_{\text{optimization error}}$
$+$ $\underbrace{\sqrt{\tfrac{\log(n/\delta)}{m}}}_{\text{approximation error}}+\underbrace{\tfrac{d+\log(m/\delta^{\prime})}{\sqrt{m}}}_{\text{linearization error}}$
with an appropriate step size, projection radius, and with probability at least $1-\delta-\delta^{\prime}$ over the random initialization (introduced in Section 5.2).
For ProjSGD, the same bound holds for $\mathbb{E}[\widehat{\mathcal{L}}_{n}(\bar{\varphi}^{(\tau)})\,|\,\varphi^{(0)}]$ with the same rates (cf. Proposition 1). We highlight the following features:
•  Non-asymptotic in time and width. The training loss decays at the canonical $1/\sqrt{\tau}$ optimization rate and $1/\sqrt{m}$ width rate.
•  Logarithmic in sample size. The dependence on the sample size enters only via $\sqrt{\log(n/\delta)/m}$, i.e., logarithmically in $n$.
•  No explicit dependence on sequence length $T$. The displayed rate has no $T$ dependence thanks to the stable structure of the attention gradients.
5 TRANSFORMER NTK ANALYSIS
In this section, for the Transformer architecture introduced in Section 3.2, we analyze its gradients in Section 5.1, introduce the initialization in Section 5.2, derive the Transformer NTK in Section 5.3, and introduce the concept of transportation mappings in Section 5.4. These tools define the function class against which our main results are established in Section 5.5. Proofs of the stated results are provided in the appendices as part of the supplementary material.
5.1 Transformer Gradient Analysis
Lemma 1 (Gradients).
Fix $i\in[m]$ and write $a_{i}:=a(X;W_{i})$. Then
$\frac{\partial f(X;\varphi)}{\partial c_{i}}$ $=m^{-1/2}\,h(X;\theta_{i}),$ (5)
$\nabla_{U_{i}}f(X;\varphi)$ $=m^{-1/2}\,c_{i}\,\sigma^{\prime}\!\big(U_{i}^{\top}a_{i}\big)\,a_{i},$ (6)
$\nabla_{W_{i}}f(X;\varphi)$ $=m^{-1/2}\,c_{i}\,\sigma^{\prime}\!\big(U_{i}^{\top}a_{i}\big)\,\big(M_{i}U_{i}\big)\,q_{X}^{\top},$ (7)
where
$M_{i}\;:=\;M(X;W_{i})\;=\;X\,J_{s}\!\big(X^{\top}W_{i}q_{X}\big)\,X^{\top},$ (8)
and $J_{s}(z)\;=\;\operatorname{diag}\!\big(\sigma_{s}(z)\big)-\sigma_{s}(z)\sigma_{s}(z)^{\top}$ is the Jacobian matrix of softmax.
Remark 4.
With $\alpha=\sigma_{s}(X^{\top}W_{i}q_{X})\in\Delta_{T}$ and $\mu=\sum_{t}\alpha_{t}X_{t}$, notice that $M(X;W_{i})=\sum_{t=1}^{T}\alpha_{t}X_{t}X_{t}^{\top}-\mu\mu^{\top}$ is a softmax‑weighted covariance that appears in (7) as a result of the softmax Jacobian matrix. Since $\mathrm{tr}\,M(X;W_{i})\leq 1$ for all $X\in\mathcal{X}$, we obtain $\|M(X;W_{i})\|_{F}\leq 1$ uniformly in $T$. As a consequence of this key property, we observe that $\|\nabla_{W_{i}}f(X;\varphi)\|_{F}$ is independent of $T$. In contrast, for RNNs the analogous gradient contains an explicit sum over $T$ recurrent steps, leading to constants that can grow with $T$; see (Cayci and Eryilmaz, 2024a, Prop. 3.1). This structural difference underlies our TT‑independent optimization bounds in Section 5.5.
Remark 5.
Similarly, we have $a(X;W_{i})=\sum_{t=1}^{T}\alpha_{t}X_{t}$, which is a convex combination of the tokens $\{X_{t}\}_{t=1}^{T}$. Hence, by the triangle inequality, $\|a(X;W_{i})\|_{2}\;\leq\;\sum_{t=1}^{T}\alpha_{t}\|X_{t}\|_{2}\;\leq\;\sum_{t=1}^{T}\alpha_{t}\;=\;1$. This allows us to bound $\|\nabla_{U_{i}}f(X;\varphi)\|_{2}$, and together with the previous remark, these bounds yield the continuity and smoothness results stated in Lemma 2.
5.2 Initialization
As noted in previous work (Wu et al., 2023; Cayci and Eryilmaz, 2024a; Lee et al., 2019; Jacot et al., 2018; Chizat et al., 2019), initialization plays an important role in the performance of gradient descent in the kernel regime. Following (Bai and Lee, 2020; Cayci and Eryilmaz, 2024a; Chizat et al., 2019), we adopt a symmetric random initialization that centers the model at zero output, which will be important in the convergence analysis. Assume $m$ is even. For $i\in[m/2]$ initialize
$W_{i}^{(0)}$ $\in\mathbb{R}^{d\times d}\ \text{with i.i.d.\ entries }\mathcal{N}(0,1),$
$U_{i}^{(0)}$ $\sim\mathcal{N}(0,I_{d}),\qquad c_{i}^{(0)}\sim\mathrm{Rad}(1).$
and set
$W^{(0)}_{i+m/2}=W^{(0)}_{i},\quad U^{(0)}_{i+m/2}=U^{(0)}_{i},\quad c^{(0)}_{i+m/2}=-\,c_{i}^{(0)},$
with all draws mutually independent. We will denote this random initialization by $\varphi^{(0)}$.
Remark 6.
Writing $h_{i}(X)=h(X;\theta_{i}^{(0)})$, note that the pairing above gives $h_{i+m/2}(X)=h_{i}(X)$, and therefore $f\bigl(X;\varphi^{(0)}\bigr)=0$ for all $X\in\mathcal{X}$. This fact will be critical to achieve the error bounds in Lemmas 3 and 5.
Because $U_{i}^{(0)}$ is a sequence of sub-Gaussian vectors, the running maximum of its norm stays bounded with high probability, which allows us to bound the norm of the attention weight gradient in (7) with high probability. For this, we define the following event.
Definition 1.
For $\delta^{\prime}\in(0,1)$, define
$\mathcal{E}_{U,m}(\delta^{\prime}):=\Big\{\,\max_{i\in[m]}\|U_{i}^{(0)}\|_{2}\;\leq\;\sqrt{d}\;+\;\sqrt{2\log(m/2\delta^{\prime})}\,\Big\}.$
By Gaussian concentration for the $1$-Lipschitz map $u\mapsto\|u\|_{2}$ (see (Boucheron et al., 2013, Theorem 5.6)) and a union bound over $i\in[m/2]$, we obtain $\mathbb{P}\big(\mathcal{E}_{U,m}(\delta^{\prime})\big)\;\geq\;1-\delta^{\prime}.$
Throughout the paper, we will work in the following parameter set centered around the initialization $\varphi^{(0)}$.
Definition 2.
Let $\rho=(\rho_{c},\rho_{u},\rho_{w})\in\mathbb{R}_{+}^{3}$. For $i\in[m]$ define
$\mathcal{C}_{\rho_{c},i}$ $:=\{\,c\in\mathbb{R}:|c-c_{i}^{(0)}|\leq\rho_{c}/\sqrt{m}\,\},$
$\mathcal{U}_{\rho_{u},i}$ $:=\{\,U\in\mathbb{R}^{d}:\|U-U_{i}^{(0)}\|_{2}\leq\rho_{u}/\sqrt{m}\,\},$
$\mathcal{W}_{\rho_{w},i}$ $:=\{\,W\in\mathbb{R}^{d\times d}:\|W-W_{i}^{(0)}\|_{F}\leq\rho_{w}/\sqrt{m}\,\},$
and the product set
$\Omega_{\rho,i}\;:=\;\mathcal{C}_{\rho_{c},i}\times\mathcal{U}_{\rho_{u},i}\times\mathcal{W}_{\rho_{w},i},\qquad\Omega_{\rho}\;:=\;\prod_{i=1}^{m}\Omega_{\rho,i}.$
Remark 7.
Notice that on $\mathcal{E}_{U,m}(\delta^{\prime})$ we also have for all $i\in[m]$ and all $U_{i}\in\mathcal{U}_{\rho_{u},i}$
$\|U_{i}\|_{2}\leq\sqrt{d}+\sqrt{2\log(m/2\delta^{\prime})}+\frac{\rho_{u}}{\sqrt{m}}=:B_{U,m}(\delta^{\prime}).$
This allows us to establish the local Lipschitz continuity and smoothness, which in turn enable us to control the optimization error in Theorem 1.
Lemma 2.
On $\mathcal{E}_{U,m}(\delta^{\prime})$, the mapping $\theta_{i}\mapsto h(X;\theta_{i})$ is $L_{1,m}(\delta^{\prime})$-Lipschitz and $L_{2,m}(\delta^{\prime})$-smooth in $\Omega_{\rho}$, where $L_{1,m}(\delta^{\prime}):=\sigma_{1}\sqrt{1+B_{U,m}(\delta^{\prime})^{2}},$ and
$L_{2,m}(\delta^{\prime}):=\sigma_{2}(1+B_{U,m}(\delta^{\prime})^{2})+8\sigma_{1}\sqrt{1+B_{U,m}(\delta^{\prime})^{2}}.$
Next, we define the linear model, which will be useful in the analysis.
Definition 3 (Linear model).
Define the function
$f_{\mathrm{lin}}(X;\varphi):=f(X;\varphi^{(0)})\;+\;\big\langle\nabla_{\varphi}f(X;\varphi^{(0)}),\,\varphi-\varphi^{(0)}\big\rangle,$
which is the first-order approximation of $f$ at initialization. Thanks to symmetric initialization, we get
$f_{\mathrm{lin}}(X;\varphi)=\big\langle\nabla_{\varphi}f(X;\varphi^{(0)}),\,\varphi-\varphi^{(0)}\big\rangle.$
Near initialization, higher-order errors vanish, so the linear model closely approximates the original model. Indeed, inside the neighborhood $\Omega_{\rho}$, we get the following uniform linearization error bound.
Lemma 3 (Linearization error).
Assume $\varphi\in\Omega_{\rho}$ and define $\varepsilon_{\mathrm{lin}}:=\sup_{j\in[n]}\big|f(X^{(j)};\varphi)-f_{\mathrm{lin}}(X^{(j)};\varphi)\big|$. Conditioned on the event $\mathcal{E}_{U,m}(\delta^{\prime})$, we have
$\varepsilon_{\mathrm{lin}}\leq\tfrac{1}{\sqrt{m}}\Bigl(L_{1,m}(\delta^{\prime})\rho_{c}\sqrt{\rho_{w}^{2}+\rho_{u}^{2}}+L_{2,m}(\delta^{\prime})(\rho_{w}^{2}+\rho_{u}^{2})\Bigr),$
where the upper bound is denoted as $B_{\mathrm{lin},m}(\delta^{\prime}).$
5.3 Transformer NTK
Let $\varphi^{(0)}$ be the random initialization. The NTK is defined as
$K(X,X^{\prime})\coloneqq\lim_{m\to\infty}\Big\langle\nabla_{\varphi}f\!\big(X;\varphi^{(0)}\big),\,\nabla_{\varphi}f\!\big(X^{\prime};\varphi^{(0)}\big)\Big\rangle.$
Viewing $\nabla_{\varphi}f(X;\varphi^{(0)})$ as a feature map for the input $X$, $K$ is the kernel induced by these features in the infinite-width limit. In this regime, gradient descent on $f$ linearizes around $\varphi^{(0)}$ and corresponds to kernel regression with kernel $K$.
Because $\nabla_{\varphi}f(X;\varphi^{(0)})$ is block-separable across $(c,U,W)$, we investigate the features coming from these parameters separately. Using the gradient identities in Lemma 1, the Transformer NTK decomposes into the following blocks.
Lemma 4 (NTK decomposition).
Let $a=a(X;W)$, $a^{\prime}=a(X^{\prime};W)$, $M=X\,J_{s}(X^{\top}Wq_{X})\,X^{\top}$, and $M^{\prime}=X^{\prime}\,J_{s}(X^{\prime\top}Wq_{X^{\prime}})\,X^{\prime\top}$. Then, with expectation taken over the random initialization $(c,U,\operatorname{vec}(W))\sim\varphi^{(0)}_{1}$,
$K_{c}(X,X^{\prime})$ $=\mathbb{E}\!\left[\,\sigma(U^{\top}a)\,\sigma(U^{\top}a^{\prime})\,\right],$
$K_{u}(X,X^{\prime})$ $=\mathbb{E}\!\left[\,\sigma^{\prime}(U^{\top}a)\,\sigma^{\prime}(U^{\top}a^{\prime})\,\langle a,a^{\prime}\rangle\,\right],$
$K_{w}(X,X^{\prime})$ $=\langle q_{X},q_{X^{\prime}}\rangle\mathbb{E}\!\left[\sigma^{\prime}(U^{\top}a)\sigma^{\prime}(U^{\top}a^{\prime})U^{\top}MM^{\prime}U\right]$
and the NTK decomposes as $K=K_{c}+K_{u}+K_{w}$.
Remark 8.
Unlike the NTK analysis in (Wu et al., 2023, Lemma 1), here the attention‑weight component $K_{w}$ is nonzero. This is a consequence of not using the degenerate $m^{-1}$ scaling and keeping a random attention weight in (2), which preserves attention features in the limit. Moreover, with multiple independent heads, the empirical NTK converges to the deterministic kernel $K$ above, solving the limit kernel problem of a single-head attention mentioned in (Hron et al., 2020).
5.4 Transportation Mappings
Let $\varphi_{0}:=(c,U,\operatorname{vec}(W))\sim\varphi^{(0)}_{1}$. It can be seen that the NTK in Lemma 4 is a Mercer kernel constructed with the following random feature maps:
$\phi_{c}(X)$ $:=\sigma\!\big(U^{\top}a(X;W)\big)\in\mathbb{R},$
$\phi_{u}(X)$ $:=\sigma^{\prime}\!\big(U^{\top}a(X;W)\big)\,a(X;W)\in\mathbb{R}^{d},$
$\phi_{w}(X)$ $:=\sigma^{\prime}\!\big(U^{\top}a(X;W)\big)\,(M(X;W)U)\,q_{X}^{\top}\in\mathbb{R}^{d\times d}.$
In order to characterize the Reproducing Kernel Hilbert Space (RKHS) associated with the NTK, we will use the concept of transportation mappings (Ji et al., 2020).
Define a measurable $v:\mathbb{R}^{d^{2}+d+1}\to\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R}^{d\times d}$ with $v(\varphi_{0})=(v_{c}(\varphi_{0}),v_{u}(\varphi_{0}),v_{w}(\varphi_{0}))$ such that $\mathbb{E}\!\Big[\,|v_{c}(\varphi_{0})|^{2}+\|v_{u}(\varphi_{0})\|_{2}^{2}+\|v_{w}(\varphi_{0})\|_{F}^{2}\,\Big]<\infty$. Define the function class
$\mathcal{F}:=\Big\{\tilde{f}(\cdot;v):\mathbb{E}|v_{c}(\varphi_{0})|^{2}<\infty,\;\mathbb{E}\|v_{u}(\varphi_{0})\|_{2}^{2}<\infty,\\ \mathbb{E}\|v_{w}(\varphi_{0})\|_{F}^{2}<\infty\Big\},$
where
$\tilde{f}(X;v):=\mathbb{E}\!\big[\phi_{c}(X)\,v_{c}(\varphi_{0})\big]+\mathbb{E}\!\big[\langle\phi_{u}(X),v_{u}(\varphi_{0})\rangle\big]\\ +\mathbb{E}\!\big[\langle\phi_{w}(X),v_{w}(\varphi_{0})\rangle_{F}\big]$ (9)
and equip $\mathcal{F}$ with the inner product
$\langle\tilde{f}(\cdot;v),\tilde{f}(\cdot;v^{\prime})\rangle_{\mathcal{F}}:=\mathbb{E}\!\Big[v_{c}\,v_{c}^{\prime}+\langle v_{u},v_{u}^{\prime}\rangle+\langle v_{w},v_{w}^{\prime}\rangle_{F}\Big].$
Notice that for a fixed $X^{\prime}\in\mathcal{X}$, one can set $v^{\prime}(\varphi_{0}):=\left(\phi_{c}(X^{\prime}),\phi_{u}(X^{\prime}),\phi_{w}(X^{\prime})\right)$ to get
$\tilde{f}(X;v^{\prime})=\mathbb{E}\!\big[\phi_{c}(X)\phi_{c}(X^{\prime})\big]+\mathbb{E}\!\big[\langle\phi_{u}(X),\phi_{u}(X^{\prime})\rangle\big]\\ +\mathbb{E}\!\big[\langle\phi_{w}(X),\phi_{w}(X^{\prime})\rangle_{F}\big]=K(X,X^{\prime}).$
In other words, we have $K(\cdot,X^{\prime})\in\mathcal{F}$ for every $X^{\prime}\in\mathcal{X}$, and $\langle\tilde{f}(\cdot;v),K(\cdot,X^{\prime})\rangle_{\mathcal{F}}=\tilde{f}(X^{\prime};v)$. Hence, $K$ has the reproducing property on $\mathcal{F}$, which implies that the completion $\overline{\mathcal{F}}$ is the RKHS associated with the Transformer NTK $K$.
We will consider the norm-constrained functions from this class. In particular, for a given $\bar{\nu}=(\bar{\nu}_{c},\bar{\nu}_{u},\bar{\nu}_{w})\in\mathbb{R}_{+}^{3}$, the function class of interest is defined as
$\mathcal{F}_{\bar{\nu}}:=\big\{\tilde{f}(\cdot;v):\ v\in\mathcal{V}_{\bar{\nu}}\big\},$
where
$\mathcal{V}_{\bar{\nu}}$ $:=\Big\{v:\ \sup_{\varphi}|v_{c}(\varphi)|\leq\bar{\nu}_{c},\ \sup_{\varphi}\|v_{u}(\varphi)\|_{2}\leq\bar{\nu}_{u},\$
$\hskip 30.00005pt\sup_{\varphi}\|v_{w}(\varphi)\|_{F}\leq\bar{\nu}_{w}\Big\}.$
Remark 9.
Note that we have $\|\tilde{f}\|_{\mathcal{F}_{\bar{\nu}}}\leq\|\bar{\nu}\|_{2}$. This constraint enables us to control the approximation bound in Lemma 5 and gives a significant simplification in the convergence analysis of ProjGD in Theorem 1.
For any $\tilde{f}\in\mathcal{F}_{\bar{\nu}}$, the transportation mapping $v$ associated with $\tilde{f}$ “transports” the initial weights to a point at which the model $f$ more closely approximates $\tilde{f}$. Indeed, for $i\in[m]$, define the following parameters
$\tilde{c}_{i}$ $=c_{i}^{(0)}\;+\;m^{-1/2}\,v_{c}(\varphi_{i}^{(0)}),$ (10)
$\tilde{U}_{i}$ $=U_{i}^{(0)}\;+\;c_{i}^{(0)}m^{-1/2}\,v_{u}(\varphi_{i}^{(0)}),$
$\tilde{W}_{i}$ $=W_{i}^{(0)}\;+\;c_{i}^{(0)}m^{-1/2}\,v_{w}(\varphi_{i}^{(0)}),$
and $\tilde{\varphi}:=(\tilde{c}_{i},\tilde{U}_{i},\operatorname{vec}(\tilde{W}_{i}))_{i=1}^{m}$. Then we have the following approximation error bound.
Lemma 5 (Approximation error).
Let $\tilde{f}(\cdot;v)\in\mathcal{F}_{\bar{\nu}}$ and $\tilde{\varphi}$ be as in (10). Define $\varepsilon_{\mathrm{app}}:=\sup_{j\in[n]}\ \big|f_{\mathrm{lin}}(X^{(j)};\tilde{\varphi})-\tilde{f}(X^{(j)};v)\big|$. Then for any $\delta\in(0,1)$, we have
$\varepsilon_{\mathrm{app}}\;\leq\;2(\sigma_{0}\bar{\nu}_{c}+\sigma_{1}\bar{\nu}_{u}+\sigma_{1}\bar{\nu}_{w})\sqrt{\tfrac{2\log(2n/\delta)}{m}}=:B_{\mathrm{app},m}(\delta),$
with probability at least $1-\delta$. We denote this event with $\mathcal{E}_{\mathrm{app}}(\delta)$.
Remark 10.
Contrasting this result with the RNN counterpart (Cayci and Eryilmaz, 2024a, Proposition 3.5), we can see that stable attention gradient norms give an approximation bound that does not depend on $T$.
5.5 Main Results
We now present the main theorem, which formally establishes the result stated in Section 4.
Theorem 1 (ProjGD Convergence).
Assume $f^{\star}\in\mathcal{F}_{\bar{\nu}}$ and the projection radius satisfies $\rho\geq\bar{\nu}$. Run ProjGD onto $\Omega_{\rho}$ for $\tau$ steps with the step size $\eta=1/\sqrt{\tau}$. Then, conditioned on the event $\mathcal{E}_{U,m}(\delta^{\prime})\cap\mathcal{E}_{\mathrm{app}}(\delta)$, we have
$\min_{s<\tau}\ \widehat{\mathcal{L}}_{n}\!\big(\varphi^{(s)}\big)\;\lesssim_{\bar{\sigma},\bar{\nu},\rho}\!\;\frac{L_{1,m}(\delta^{\prime})^{4}}{\sqrt{\tau}}\\ +L_{1,m}(\delta^{\prime})(B_{\mathrm{app},m}(\delta)+B_{\mathrm{lin},m}(\delta^{\prime})+\varepsilon_{\mathrm{CoF}}),$
and for the average iterate $\bar{\varphi}^{(\tau)}:=\frac{1}{\tau}\sum_{s=0}^{\tau-1}\varphi^{(s)}$,
$\widehat{\mathcal{L}}_{n}\!\big(\bar{\varphi}^{(\tau)}\big)\;\lesssim_{\bar{\sigma},\bar{\nu},\rho}\!\;\frac{L_{1,m}(\delta^{\prime})^{4}}{\sqrt{\tau}}\\ +L_{1,m}(\delta^{\prime})(B_{\mathrm{app},m}(\delta)+B_{\mathrm{lin},m}(\delta^{\prime})+\varepsilon_{\mathrm{CoF}})+B_{\mathrm{lin},m}(\delta^{\prime})^{2},$
where $B_{\mathrm{lin},m}(\delta^{\prime})$ and $B_{\mathrm{app},m}(\delta)$ are defined as in Lemmas 3 and 5, and
$\varepsilon_{\mathrm{CoF}}\;\lesssim_{\bar{\sigma},\bar{\nu},\rho}\;\frac{L_{1,m}(\delta^{\prime})+L_{2,m}(\delta^{\prime})}{\sqrt{m}}.$
We can state a similar result for the convergence of ProjSGD as well.
Proposition 1 (ProjSGD Convergence).
Under the same conditions as Theorem 1, we have
$\mathbb{E}\left[\min_{s<\tau}\widehat{\mathcal{L}}_{n}(\varphi^{(s)})\,|\,\varphi^{(0)}\right]\;\lesssim_{\bar{\sigma},\bar{\nu},\rho}\!\;\frac{L_{1,m}(\delta^{\prime})^{4}}{\sqrt{\tau}}\\ +L_{1,m}(\delta^{\prime})(B_{\mathrm{app},m}(\delta)+B_{\mathrm{lin},m}(\delta^{\prime})+\varepsilon_{\mathrm{CoF}}),$
and for the average iterate,
$\mathbb{E}\left[\widehat{\mathcal{L}}_{n}\!\big(\bar{\varphi}^{(\tau)}\big)\,|\,\varphi^{(0)}\right]\!\;\lesssim_{\bar{\sigma},\bar{\nu},\rho}\!\;\frac{L_{1,m}(\delta^{\prime})^{4}}{\sqrt{\tau}}\\ +L_{1,m}(\delta^{\prime})(B_{\mathrm{app},m}(\delta)+B_{\mathrm{lin},m}(\delta^{\prime})+\varepsilon_{\mathrm{CoF}})+B_{\mathrm{lin},m}(\delta^{\prime})^{2}.$
6 EXPERIMENTAL RESULTS
In Section 6.1, we validate the scaling laws predicted by Lemma 3, Lemma 5, and Theorem 1. In Section 6.2, we compare IndRNN and our Transformer model on an autoregressive forecasting task, highlighting differences in handling long-term dependencies and the accompanying memory trade-offs.
6.1 Teacher–Student Scaling-Law Validation
We assess the three components of our theory—linearization, approximation, and optimization—with controlled teacher–student experiments using the same shallow Transformer as in Section 3.2 with $q_{X}=X_{T}$. Throughout, inputs are i.i.d. $X\sim\mathcal{N}(0,I)$ with token-wise clipping to keep $\max_{t}\|X_{t}\|_{2}\leq 1$.
Teacher.
To realize a ground-truth $f^{\star}\in\mathcal{F}_{\bar{\nu}}$ as in (9) we define a sup-norm-bounded transportation mapping $v\in\mathcal{V}_{\bar{\nu}}$ as $v(\varphi)=\frac{1}{R}\sum_{r=1}^{R}\,v^{(r)}(\varphi)$, where
$v_{c}^{(r)}(\varphi)$ $=\tfrac{\nu_{c}}{\sigma_{0}}\,\phi_{c}(X_{R}^{(r)}),\quad v_{u}^{(r)}(\varphi)=\tfrac{\nu_{u}}{\sigma_{1}}\,\phi_{u}(X_{R}^{(r)}),$
$v_{w}^{(r)}(\varphi)$ $=\operatorname{clip}_{F}\!\big(\phi_{w}(X_{R}^{(r)}),\,\nu_{w}\big),$
and the anchor samples $X_{R}^{(r)}$ are drawn from the same distribution as inputs once and kept fixed. For a large $\tilde{m}\in\mathbb{N}$, we sample an i.i.d. pool $\{\varphi_{i}^{(0)}\}_{i=1}^{\tilde{m}}$ and generate the labels via a Monte Carlo approximation of (9), i.e.,
$\tilde{y}^{(j)}=\frac{1}{\tilde{m}}\sum_{i=1}^{\tilde{m}}\!\big[\phi_{c}(X^{(j)})\,v_{c}(\varphi_{i}^{(0)})+\langle\phi_{u}(X^{(j)}),v_{u}(\varphi_{i}^{(0)})\rangle\\ +\langle\phi_{w}(X^{(j)}),v_{w}(\varphi_{i}^{(0)})\rangle_{F}\big],$
so the teacher lies in the RKHS induced by the model’s NTK, matching the realizability assumption used in the proofs. All experiments are conducted with $\tilde{m}=8192$, $d=8$, $T=16$, $n=5000$, $\bar{\nu}=(3,3,3)$, and $\tanh$ activation.
Students and metrics.
Students are initialized with the symmetric initialization from Section 5.2 so that $f(\cdot;\varphi^{(0)})\equiv 0$. For each width $m\in\{8,16,64,128,256\}$, we record:
•  Linearization error. We select parameters on the boundary of the neighborhood $\Omega_{\rho}$ and compute $\sup_{j\in[n]}\big|f(X^{(j)};\varphi)-f_{\mathrm{lin}}(X^{(j)};\varphi)\big|$, which should scale like $m^{-1/2}$ per Lemma 3.
•  Approximation error. We set the parameters $\tilde{\varphi}$ as in (10) and measure $\sup_{j\in[n]}\ \big|f_{\mathrm{lin}}(X^{(j)};\tilde{\varphi})-\tilde{y}^{(j)}\big|$, which is expected to scale like $m^{-1/2}$ from Lemma 5.
•  Minimum training loss. We train a student Transformer with $\tau=4000$ steps, a constant step size $\eta=1/\sqrt{\tau}$, and projection radii $\rho=\bar{\nu}$. Minimum training loss $\min_{t<\tau}\widehat{\mathcal{L}}_{n}(\varphi^{(t)})$ versus $m$ is supposed to scale like $m^{-1/2}$ from Theorem 1.
(a) Linearization error
(b) Approximation error
(c) Min training loss
Figure 1: Width scaling across three metrics (log–log). Shaded bands show mean $\pm$95% CI across seeds. Fitted slopes are consistent with the predicted $m^{-1/2}$ decay.
We average curves over $10$ seeds, show $\pm$95% confidence interval (CI) around the mean, and fit a straight line to the log–log curves to estimate the scaling exponent. It can be seen from Figure 1 that all three quantities—linearization error, approximation error, and minimum training loss—decrease monotonically with width and exhibit a clear power-law dependence on $m$. The fitted slopes are in close agreement with the predicted $m^{-1/2}$ rate. Medians closely track the mean and the CIs remain small across widths, indicating that the observed scaling is robust to random initialization.
6.2 IndRNN vs. Transformer Architectures
To compare how a recursive model (IndRNN (Cayci and Eryilmaz, 2024a)) and our Transformer cope with long-term dependencies and training stability, we use a controlled autoregressive AR($L$) task where the lag $L$ sets the long-term dependency of the problem. Theory for diagonal RNNs predicts that depending on a Lipschitz cutoff, gradients can grow exponentially due to repeated multiplication over a horizon of $T$ steps (Cayci and Eryilmaz, 2024a), whereas our Transformer’s bound has no explicit $T$-dependence (cf. Lemma 2). Each input sequence $X\!\in\!\mathbb{R}^{1\times T}$ is constructed as a window from a univariate process
$X_{t}=\alpha\,X_{t-L}+\varepsilon_{t},\qquad\varepsilon_{t}\sim\mathcal{N}(0,\sigma_{\varepsilon}^{2}),$
and the label is $y=X_{T+1}$. To the input we append fixed positional channels ($d_{\text{pos}}{=}8$ sinusoidal pairs) so that the Transformer model has positional knowledge. We fix n=5,000n=5{,}000, $\alpha=0.9$, $\sigma_{\varepsilon}^{2}=0.1$, and vary the lag $L$ to toggle short/long temporal dependency inside the window. For Transformer, we set $T=L+1$ so that the lag is inside the context window, which makes Transformer’s memory footprint $\mathcal{O}(L)$ compared to IndRNN’s $\mathcal{O}(1)$ footprint. Both models are initialized with a symmetric initialization and are trained with gradient descent for $\tau=2{,}000$ steps with step size $\eta=1/\sqrt{\tau}$ and no projection to observe the gradient norm dynamics. We set $m=64$ and use $\tanh$. The results are averaged over $20$ seeds.
Figure 2: Performance comparison of IndRNN and Transformer architectures with varying time lag $L$.
In Figure 2, we contrast the function Jacobian norms and the minimum validation loss performances of the IndRNN and Transformer architectures for a range of lag values. In particular, the upper plot compares the maximum Jacobian norm occurred in training with respect to the diagonal recurrent weights for IndRNN and attention weights for Transformer. We observe that maximum IndRNN norms are large with higher fluctuations, which confirms the theory established in (Cayci and Eryilmaz, 2024a); whereas Transformer norms consistently take smaller values as predicted by our theory. The lower plot compares the minimum validation loss occurred in training for both architectures. It can be seen that when the problem has a short-term dependency (low $L$), IndRNN achieves better results with degrading performance as $L$ increases; on the other hand, Transformer gives a consistent performance that does not degrade with $L$, at the expense of its growing memory footprint. As such, this highlights the performance-memory trade-off between the two architectures that is anticipated by the theory.
7 CONCLUSION
We performed a finite-time NTK analysis of projected gradient methods for multi-head shallow Transformers. We established nonasymptotic convergence bounds with explicit dependencies on the training period, network width, ambient dimension, and the number of samples. To the best of our knowledge, this is the first work that establishes near-optimality of first-order methods without massive overparameterization for Transformers. Interesting future directions that can build on this work include the study of deep Transformer architectures and the extension of the setting to sequence-to-sequence learning.
Supplementary Material
Organization.
In Appendix A, we restate and prove the results in the main text, and in Appendix B, we give further experimental details along with the anonymous code repository to reproduce the experiments.
Appendix A PROOFS
A.1 Useful facts
Fact 1 (Jacobian of softmax).
For $\sigma_{s}:\mathbb{R}^{T}\to\Delta_{T}$ given by $(\sigma_{s}(z))_{t}=\exp(z_{t})/\sum_{s=1}^{T}\exp(z_{s})$, the Jacobian equals
$J_{s}(z)\;=\;\operatorname{diag}\!\big(\sigma_{s}(z)\big)\;-\;\sigma_{s}(z)\,\sigma_{s}(z)^{\top}.$
It is symmetric, satisfies $J_{s}(z)\mathbf{1}=0$, and obeys the operator bound $\|J_{s}(z)\|_{\infty\to 1}\leq 2$. Here, $\mathbf{1}$ is the all‑ones vector.
Proof.
$\partial(\sigma_{s})_{i}/\partial z_{j}=\sigma_{s,i}(\delta_{ij}-\sigma_{s,j})$ gives the formula and symmetry. Row sums vanish since $\sum_{j}\partial(\sigma_{s})_{i}/\partial z_{j}=0$, hence $J_{s}(z)\mathbf{1}=0$. For the $\infty\!\to\!1$ bound, if $\|u\|_{\infty}\leq 1$ then $\|J_{s}(z)u\|_{1}=\|\operatorname{diag}(\alpha)u-\alpha(\alpha^{\top}u)\|_{1}\leq\sum_{i}\alpha_{i}|u_{i}|+|\alpha^{\top}u|\,\|\alpha\|_{1}\leq 1+1=2$ with $\alpha=\sigma_{s}(z)$. ∎
Fact 2 (Bound for $a(X;W)$).
Fix $X\in\mathcal{X}$. With $\alpha=\sigma_{s}(X^{\top}Wq_{X})\in\Delta_{T}$,
$a(X;W)\;=\;X\alpha\;=\;\sum_{t=1}^{T}\alpha_{t}\,X_{t}\quad\Rightarrow\quad\|a(X;W)\|_{2}\;\leq\;\sum_{t=1}^{T}\alpha_{t}\|X_{t}\|_{2}\;\leq\;\sum_{t=1}^{T}\alpha_{t}\;=\;1.$
Fact 3 (Bound for $M(X;W)$).
Fix $X\in\mathcal{X}$. Let $\alpha=\sigma_{s}(X^{\top}Wq_{X})$ and $\mu=X\alpha$. Then
$M(X;W)\;=\;X\big(\operatorname{diag}(\alpha)-\alpha\alpha^{\top}\big)X^{\top}\;=\;\sum_{t=1}^{T}\alpha_{t}\,(X_{t}-\mu)(X_{t}-\mu)^{\top}\;\succeq\;0.$
Consequently,
$M\;\preceq\;\sum_{t=1}^{T}\alpha_{t}\,X_{t}X_{t}^{\top}\quad\text{and}\quad\|M\|_{F}\;\leq\;\mathrm{tr}(M)\;\leq\;\sum_{t=1}^{T}\alpha_{t}\|X_{t}\|_{2}^{2}\;\leq\;1.$
Fact 4 (Gaussian norm concentration).
If $G\sim\mathcal{N}(0,I_{d})$, then
$\mathbb{P}\!\left(\big|\|G\|_{2}-\mathbb{E}\|G\|_{2}\big|\geq t\right)\;\leq\;2e^{-t^{2}/2}\quad\text{for all }t>0,\qquad\mathbb{E}\|G\|_{2}\;\leq\;\sqrt{d}.$
In particular, for independent $G_{1},\dots,G_{m}\sim\mathcal{N}(0,I_{d})$,
$\mathbb{P}\!\left(\max_{1\leq i\leq m}\|G_{i}\|_{2}\;\leq\;\sqrt{d}+\sqrt{2\log(m/\delta)}\right)\;\geq\;1-\delta.$
Proof.
By Lipschitz Gaussian concentration (Boucheron et al., 2013, Theorem 5.6) applied to $u\mapsto\|u\|_{2}$, we get the tail bound; the expectation bound follows from $\mathbb{E}\|G\|_{2}\leq\sqrt{\mathbb{E}\|G\|_{2}^{2}}=\sqrt{d}$. The union bound over $m$ copies yields the shown high‑probability bound. ∎
A.2 Proof of Lemma 1
Lemma 1.
Fix $i\in[m]$ and write $a_{i}:=a(X;W_{i})$. Then
$\frac{\partial f(X;\varphi)}{\partial c_{i}}$ $=m^{-1/2}\,h(X;\theta_{i}),$ (11)
$\nabla_{U_{i}}f(X;\varphi)$ $=m^{-1/2}\,c_{i}\,\sigma^{\prime}\!\big(U_{i}^{\top}a_{i}\big)\,a_{i},$ (12)
$\nabla_{W_{i}}f(X;\varphi)$ $=m^{-1/2}\,c_{i}\,\sigma^{\prime}\!\big(U_{i}^{\top}a_{i}\big)\,\big(M_{i}U_{i}\big)\,q_{X}^{\top},$ (13)
where
$M_{i}\;:=\;M(X;W_{i})\;=\;X\,J_{s}\!\big(X^{\top}W_{i}q_{X}\big)\,X^{\top},$ (14)
and $J_{s}(z)\;=\;\operatorname{diag}\!\big(\sigma_{s}(z)\big)-\sigma_{s}(z)\sigma_{s}(z)^{\top}$ is the Jacobian matrix of softmax.
Proof.
Write $\alpha_{i}\coloneqq\sigma_{s}\!\big(X^{\top}W_{i}q_{X}\big)\in\Delta_{T}$, $a_{i}=X\alpha_{i}\in\mathbb{R}^{d}$, and $z_{i}\coloneqq X^{\top}W_{i}q_{X}\in\mathbb{R}^{T}$.
Derivative w.r.t. $c_{i}$.
From $f(X;\varphi)=\tfrac{1}{\sqrt{m}}\sum_{j=1}^{m}c_{j}h(X;\theta_{j})$, $\partial f/\partial c_{i}=m^{-1/2}\,h(X;\theta_{i})$.
Gradient w.r.t. $U_{i}$.
Since $h(X;\theta_{i})=\sigma(U_{i}^{\top}a_{i})$,
$\nabla_{U_{i}}h(X;\theta_{i})\;=\;\sigma^{\prime}\!\big(U_{i}^{\top}a_{i}\big)\,a_{i},\qquad\nabla_{U_{i}}f\;=\;m^{-1/2}\,c_{i}\,\nabla_{U_{i}}h\;=\;m^{-1/2}\,c_{i}\,\sigma^{\prime}\!\big(U_{i}^{\top}a_{i}\big)\,a_{i}.$
Gradient w.r.t. $W_{i}$.
We compute via differentials. First, by the chain rule for softmax,
$d\alpha_{i}\;=\;J_{s}(z_{i})\,dz_{i},\qquad dz_{i}\;=\;X^{\top}\,dW_{i}\,q_{X},$
and hence
$da_{i}\;=\;X\,d\alpha_{i}\;=\;X\,J_{s}(z_{i})\,X^{\top}\,dW_{i}\,q_{X}\;=\;M_{i}\,dW_{i}\,q_{X}.$
Next,
$dh(X;\theta_{i})\;=\;\sigma^{\prime}\!\big(U_{i}^{\top}a_{i}\big)\,U_{i}^{\top}\,da_{i}\;=\;\sigma^{\prime}\!\big(U_{i}^{\top}a_{i}\big)\,U_{i}^{\top}\,M_{i}\,dW_{i}\,q_{X}.$
Therefore
$df\;=\;m^{-1/2}\,c_{i}\,dh\;=\;m^{-1/2}\,c_{i}\,\sigma^{\prime}\!\big(U_{i}^{\top}a_{i}\big)\,U_{i}^{\top}M_{i}\,dW_{i}\,q_{X}.$
Using $\operatorname{tr}(A^{\top}B)=\langle A,B\rangle_{F}$ and cyclicity of trace,
$U_{i}^{\top}M_{i}\,dW_{i}\,q_{X}\;=\;\operatorname{tr}\!\big(q_{X}\,U_{i}^{\top}M_{i}\,dW_{i}\big)\;=\;\operatorname{tr}\!\big((M_{i}U_{i}q_{X}^{\top})^{\top}dW_{i}\big),$
where we used that $M_{i}$ is symmetric. Identifying the coefficient of $dW_{i}$ in the Frobenius inner product yields
$\nabla_{W_{i}}f(X;\varphi)\;=\;m^{-1/2}\,c_{i}\,\sigma^{\prime}\!\big(U_{i}^{\top}a_{i}\big)\,(M_{i}U_{i})\,q_{X}^{\top},$
as claimed. ∎
A.3 Proof of Lemma 2
Lemma 2.
On $\mathcal{E}_{U,m}(\delta^{\prime})$, the mapping $\theta_{i}\mapsto h(X;\theta_{i})$ is $L_{1,m}(\delta^{\prime})$-Lipschitz and $L_{2,m}(\delta^{\prime})$-smooth in $\Omega_{\rho}$, where
$L_{1,m}(\delta^{\prime}):=\sigma_{1}\sqrt{1+B_{U,m}(\delta^{\prime})^{2}},$
and
$L_{2,m}(\delta^{\prime}):=\sigma_{2}(1+B_{U,m}(\delta^{\prime})^{2})+8\sigma_{1}\sqrt{1+B_{U,m}(\delta^{\prime})^{2}},$
where $B_{U,m}(\delta^{\prime})$ is defined as in Remark 7.
Proof.
Throughout, $D[\cdot]$ denotes the Fréchet derivative. For a scalar map $g:\mathbb{R}^{d\times d}\to\mathbb{R}$ the Fréchet derivative is $Dg(W)[\Delta]=\left.\frac{d}{dt}g(W+t\Delta)\right|_{t=0}$. By the Riesz representation theorem, there is a unique $\nabla_{W}g(W)$ with $Dg(W)[\Delta]=\langle\nabla_{W}g(W),\Delta\rangle_{F}$, hence $\|\nabla_{W}g(W)\|_{F}=\sup_{\|\Delta\|_{F}=1}|Dg(W)[\Delta]|.$ For a matrix-valued operator $G$, we measure $DG(W)$ by the induced norm $\|DG(W)\|_{F\to F}:=\sup_{\|\Delta\|_{F}=1}\|DG(W)[\Delta]\|_{F}$.
Fix $i\in[m]$ and write
$\alpha_{i}\;=\;\sigma_{s}\!\big(X^{\top}W_{i}q_{X}\big)\in\Delta_{T},\qquad a_{i}\;=\;a(X;W_{i})\;=\;X\alpha_{i},\qquad z_{i}\;=\;U_{i}^{\top}a_{i},$
and
$M_{i}\;=\;X\,J_{s}\!\big(X^{\top}W_{i}q_{X}\big)\,X^{\top},$
where $J_{s}$ is the softmax Jacobian.
(a) Local Lipschitz continuity.
From Lemma 1,
$\nabla_{U_{i}}h(X;\theta_{i})=\sigma^{\prime}(z_{i})\,a_{i},\qquad\nabla_{W_{i}}h(X;\theta_{i})=\sigma^{\prime}(z_{i})\,(M_{i}U_{i})\,q_{X}^{\top}.$
Therefore
$\big\|\nabla_{U_{i}}h(X;\theta_{i})\big\|_{2}\leq\sigma_{1}\|a_{i}\|_{2}\leq\sigma_{1},\qquad\big\|\nabla_{W_{i}}h(X;\theta_{i})\big\|_{F}\leq\sigma_{1}\|M_{i}\|_{F}\|U_{i}\|_{2}\|q_{X}\|_{2}\leq\sigma_{1}\|U_{i}\|_{2}.$
Collecting the two blocks,
$\big\|\nabla_{\theta_{i}}h(X;\theta_{i})\big\|_{2}^{2}=\big\|\nabla_{U_{i}}h\big\|_{2}^{2}+\big\|\nabla_{W_{i}}h\big\|_{F}^{2}\;\leq\;\sigma_{1}^{2}\,(1+\|U_{i}\|_{2}^{2}).$
Hence, on $\mathcal{E}_{U,m}(\delta^{\prime})$, the mapping $\theta_{i}\mapsto h(X;\theta_{i})$ is $L_{1,m}(\delta^{\prime})$-Lipschitz in $\Omega_{\rho}$ with $L_{1,m}(\delta^{\prime})=\sigma_{1}\sqrt{1+B_{U,m}(\delta^{\prime})^{2}}.$
(b) Local smoothness.
Write $h_{i}:=h(X;\theta_{i})=\sigma(z_{i})$. By the chain rule,
$\nabla_{\theta_{i}}h_{i}=\sigma^{\prime}(z_{i})\,\nabla_{\theta_{i}}z_{i},\qquad\nabla_{\theta_{i}}^{2}h_{i}=\sigma^{\prime\prime}(z_{i})\,(\nabla_{\theta_{i}}z_{i})(\nabla_{\theta_{i}}z_{i})^{\top}+\sigma^{\prime}(z_{i})\,\nabla_{\theta_{i}}^{2}z_{i},$
and thus
$\big\|\nabla_{\theta_{i}}^{2}h_{i}\big\|_{F}\;\leq\;\sigma_{2}\underbrace{\big\|\nabla_{\theta_{i}}z_{i}\big\|_{2}^{2}}_{(i)}\;+\;\sigma_{1}\underbrace{\big\|\nabla_{\theta_{i}}^{2}z_{i}\big\|_{F}}_{(ii)}.$ (15)
Bounding $(i)$. We have $\nabla_{U_{i}}z_{i}=a_{i}$ and, for any direction $\Delta\in\mathbb{R}^{d\times d}$,
$Dz_{i}[\Delta]\;=\;U_{i}^{\top}M_{i}\,\Delta\,q_{X},\quad\Rightarrow\quad\|\nabla_{W_{i}}z_{i}\|_{F}\leq\|M_{i}\|_{F}\,\|U_{i}\|_{2}\,\|q_{X}\|_{2}\leq\|U_{i}\|_{2}.$
Therefore
$\big\|\nabla_{\theta_{i}}z_{i}\big\|_{2}^{2}=\|\nabla_{U_{i}}z_{i}\|_{2}^{2}+\|\nabla_{W_{i}}z_{i}\|_{F}^{2}\;\leq\;1+\|U_{i}\|_{2}^{2}.$ (16)
Bounding $(ii)$. Note $\nabla_{U_{i}}^{2}z_{i}=0$ and the $W_{i}$$U_{i}$ block of the Hessian equals the Jacobian of $a_{i}$ w.r.t. $W_{i}$:
$D\!\left[\nabla_{U_{i}}z_{i}\right][\Delta]\;=\;D[a_{i}][\Delta]\;=\;M_{i}\,\Delta\,q_{X},\qquad\Rightarrow\qquad\big\|\nabla^{2}_{U_{i},\operatorname{vec}(W_{i})}z_{i}\big\|_{F}\;\leq\;\|M_{i}\|_{F}\,\|q_{X}\|_{2}\;\leq\;1.$
For the $W_{i}$$W_{i}$ block, differentiate $\nabla_{W_{i}}z_{i}=M_{i}U_{i}q_{X}^{\top}$ once more in the direction $\Delta$:
$D\!\left[\nabla_{W_{i}}z_{i}\right][\Delta]\;=\;D[M_{i}][\Delta]\,U_{i}\,q_{X}^{\top}.$
From the covariance form of $M_{i}$, $D[M_{i}][\Delta]=\sum_{t=1}^{T}(D[\alpha_{i}][\Delta])_{t}\,(X_{t}-\mu_{i})(X_{t}-\mu_{i})^{\top},$ with $\mu_{i}:=\sum_{t=1}^{T}(\alpha_{i})_{t}X_{t}=a_{i}$ and $D[\alpha_{i}][\Delta]=J_{s}(X^{\top}W_{i}q_{X})\,X^{\top}\Delta\,q_{X}$. Hence
$\|D[M_{i}][\Delta]\|_{F}\;\leq\;\sum_{t=1}^{T}|(D\alpha_{i}[\Delta])_{t}|\,\|X_{t}-\mu_{i}\|_{2}^{2}\;\leq\;4\,\|D\alpha_{i}[\Delta]\|_{1},$
and using $\|\operatorname{diag}(p)-pp^{\top}\|_{\,\infty\to 1}\leq 2$ (Fact 1),
$\|D\alpha_{i}[\Delta]\|_{1}=\big\|(\operatorname{diag}(\alpha_{i})-\alpha_{i}\alpha_{i}^{\top})\,X^{\top}\Delta\,q_{X}\big\|_{1}\;\leq\;2\,\|X^{\top}\Delta\,q_{X}\|_{\infty}\;\leq\;2\,\|\Delta\|_{F},$
where we used $|X_{t}^{\top}\Delta q_{X}|\leq\|X_{t}\|_{2}\,\|\Delta\|_{F}\,\|q_{X}\|_{2}\leq\|\Delta\|_{F}$. Thus $\|D[M_{i}][\Delta]\|_{F}\leq 8\|\Delta\|_{F}$, and therefore
$\|D\!\left[\nabla_{W_{i}}z_{i}\right][\Delta]\|_{F}\leq\|D[M_{i}][\Delta]\|_{F}\|U_{i}\|_{2}\|q_{X}\|_{2}\leq 8\|\Delta\|_{F}\|U_{i}\|_{2},$
which implies $\big\|\nabla^{2}_{\operatorname{vec}(W_{i})}z_{i}\big\|_{F}\;\leq\;8\,\|U_{i}\|_{2}.$ Combining the $W_{i}$$U_{i}$ and $W_{i}$$W_{i}$ blocks gives
$\big\|\nabla_{\theta_{i}}^{2}z_{i}\big\|_{F}^{2}\;\leq\;2\big\|\nabla^{2}_{U_{i},\operatorname{vec}(W_{i})}z_{i}\big\|_{F}^{2}+\big\|\nabla^{2}_{\operatorname{vec}(W_{i})}z_{i}\big\|_{F}^{2}\;\leq\;2+64\,\|U_{i}\|_{2}^{2},$ (17)
hence $\|\nabla_{\theta_{i}}^{2}z_{i}\|_{F}\leq\sqrt{2+64\|U_{i}\|_{2}^{2}}\leq 8\sqrt{1+\|U_{i}\|_{2}^{2}}$.
Finally, plugging (16) and (17) into (15) yields
$\big\|\nabla_{\theta_{i}}^{2}h(X;\theta_{i})\big\|_{F}\;\leq\;\sigma_{2}\,(1+\|U_{i}\|_{2}^{2})+8\sigma_{1}\,\sqrt{1+\|U_{i}\|_{2}^{2}}.$
Therefore, on $\mathcal{E}_{U,m}(\delta^{\prime})$ and for all $U_{i}\in\mathcal{U}_{\rho_{u},i}$,
$\big\|\nabla_{\theta_{i}}h(X;\theta_{i})\big\|_{2}\leq\sigma_{1}\sqrt{1+B_{U,m}(\delta^{\prime})^{2}}\quad\text{and}\quad\big\|\nabla_{\theta_{i}}^{2}h(X;\theta_{i})\big\|_{F}\leq\sigma_{2}\!\big(1+B_{U,m}(\delta^{\prime})^{2}\big)+8\sigma_{1}\sqrt{1+B_{U,m}(\delta^{\prime})^{2}}.$
A.4 Proof of Lemma 3
Lemma 3.
Assume $\varphi\in\Omega_{\rho}$ and define $\varepsilon_{\mathrm{lin}}:=\sup_{j\in[n]}\big|f(X^{(j)};\varphi)-f_{\mathrm{lin}}(X^{(j)};\varphi)\big|$. Conditioned on the event $\mathcal{E}_{U,m}(\delta^{\prime})$, we have
$\varepsilon_{\mathrm{lin}}\leq\tfrac{1}{\sqrt{m}}\Bigl(L_{1,m}(\delta^{\prime})\rho_{c}\sqrt{\rho_{w}^{2}+\rho_{u}^{2}}+L_{2,m}(\delta^{\prime})(\rho_{w}^{2}+\rho_{u}^{2})\Bigr),$
where the upper bound is denoted as $B_{\mathrm{lin},m}(\delta^{\prime}).$
Proof.
Fix any $j\in[n]$ and write $X=X^{(j)}$. For $i\in[m]$ set $\delta c_{i}:=c_{i}-c_{i}^{(0)}$, $\delta U_{i}:=U_{i}-U_{i}^{(0)}$, $\delta W_{i}:=W_{i}-W_{i}^{(0)}$, and
$\delta\theta_{i}\;:=\;(\delta U_{i},\operatorname{vec}(\delta W_{i})),\qquad\|\delta\theta_{i}\|_{2}^{2}\;=\;\|\delta U_{i}\|_{2}^{2}+\|\delta W_{i}\|_{F}^{2}.$
We can write $f_{\mathrm{lin}}(X;\varphi)$ as
$f_{\mathrm{lin}}(X;\varphi)$ $=f\big(X;\varphi^{(0)}\big)\;+\;\sum_{i=1}^{m}\Big\langle\nabla_{c_{i}}f\big(X;\varphi^{(0)}\big),\,\delta c_{i}\Big\rangle\;+\;\sum_{i=1}^{m}\Big\langle\nabla_{\theta_{i}}f\big(X;\varphi^{(0)}\big),\,\delta\theta_{i}\Big\rangle$
$=f\big(X;\varphi^{(0)}\big)\;+\;\frac{1}{\sqrt{m}}\sum_{i=1}^{m}\Big[\,h\!\big(X;\theta_{i}^{(0)}\big)\,\delta c_{i}+c_{i}^{(0)}\,\big\langle\nabla_{\theta_{i}}h\!\big(X;\theta_{i}^{(0)}\big),\,\delta\theta_{i}\big\rangle\Big].$
A direct expansion gives the per-sample linearization error
$e_{\mathrm{lin}}(X)$ $=f(X;\varphi)-f_{\mathrm{lin}}(X;\varphi)$ (18)
$=\frac{1}{\sqrt{m}}\sum_{i=1}^{m}\delta c_{i}\,\Big(h(X;\theta_{i})-h\!\big(X;\theta_{i}^{(0)}\big)\Big)$
$\quad+\frac{1}{\sqrt{m}}\sum_{i=1}^{m}c_{i}^{(0)}\Big(h(X;\theta_{i})-h\!\big(X;\theta_{i}^{(0)}\big)-\big\langle\nabla_{\theta_{i}}h\!\big(X;\theta_{i}^{(0)}\big),\,\delta\theta_{i}\big\rangle\Big).$
where we use the fact that $f(X;\varphi^{(0)})=0$. Since $\varphi\in\Omega_{\rho}$, we have
$|\delta c_{i}|\leq\frac{\rho_{c}}{\sqrt{m}},\qquad\|\delta\theta_{i}\|_{2}\leq\frac{\sqrt{\rho_{u}^{2}+\rho_{w}^{2}}}{\sqrt{m}},\qquad\|\delta\theta_{i}\|_{2}^{2}\leq\frac{\rho_{u}^{2}+\rho_{w}^{2}}{m}.$
Using Lemma 2 and the fact that $|c_{i}^{(0)}|=1$ at initialization, on the event $\mathcal{E}_{U,m}(\delta^{\prime})$ we obtain
$|e_{\mathrm{lin}}(X)|$ $\leq\frac{1}{\sqrt{m}}\sum_{i=1}^{m}\frac{\rho_{c}}{\sqrt{m}}\cdot L_{1,m}(\delta^{\prime})\cdot\frac{\sqrt{\rho_{u}^{2}+\rho_{w}^{2}}}{\sqrt{m}}\;+\;\frac{1}{\sqrt{m}}\sum_{i=1}^{m}1\cdot\frac{L_{2,m}(\delta^{\prime})}{2}\cdot\frac{\rho_{u}^{2}+\rho_{w}^{2}}{m}$
$=\frac{1}{\sqrt{m}}\,L_{1,m}(\delta^{\prime})\,\rho_{c}\,\sqrt{\rho_{u}^{2}+\rho_{w}^{2}}\;+\;\frac{1}{\sqrt{m}}\,\frac{L_{2,m}(\delta^{\prime})}{2}\,(\rho_{u}^{2}+\rho_{w}^{2})$
$\leq\frac{1}{\sqrt{m}}\Big(L_{1,m}(\delta^{\prime})\,\rho_{c}\,\sqrt{\rho_{u}^{2}+\rho_{w}^{2}}+L_{2,m}(\delta^{\prime})\,(\rho_{u}^{2}+\rho_{w}^{2})\Big).$
This bound holds for every $X=X^{(j)}$, hence after taking a supremum over $j\in[n]$ we get
$\varepsilon_{\mathrm{lin}}\;=\;\sup_{j\in[n]}|e_{\mathrm{lin}}(X^{(j)})|\;\leq\;\frac{1}{\sqrt{m}}\Big(L_{1,m}(\delta^{\prime})\,\rho_{c}\,\sqrt{\rho_{u}^{2}+\rho_{w}^{2}}+L_{2,m}(\delta^{\prime})\,(\rho_{u}^{2}+\rho_{w}^{2})\Big)\;=:B_{\mathrm{lin},m}(\delta^{\prime}).$
A.5 Proof of Lemma 4
Lemma 4.
Let $a=a(X;W)$, $a^{\prime}=a(X^{\prime};W)$, $M=X\,J_{s}(X^{\top}Wq_{X})\,X^{\top}$, and $M^{\prime}=X^{\prime}\,J_{s}(X^{\prime\top}Wq_{X^{\prime}})\,X^{\prime\top}$. Then, with expectation taken over the random initialization $(c,U,\operatorname{vec}(W))\sim\varphi^{(0)}_{1}$,
$K_{c}(X,X^{\prime})$ $=\mathbb{E}\!\left[\,\sigma(U^{\top}a)\,\sigma(U^{\top}a^{\prime})\,\right],$
$K_{u}(X,X^{\prime})$ $=\mathbb{E}\!\left[\,\sigma^{\prime}(U^{\top}a)\,\sigma^{\prime}(U^{\top}a^{\prime})\,\langle a,a^{\prime}\rangle\,\right],$
$K_{w}(X,X^{\prime})$ $=\langle q_{X},q_{X^{\prime}}\rangle\mathbb{E}\!\left[\sigma^{\prime}(U^{\top}a)\sigma^{\prime}(U^{\top}a^{\prime})U^{\top}MM^{\prime}U\right]$
and the NTK decomposes as $K=K_{c}+K_{u}+K_{w}$.
Proof.
Define the width-$m$ neural tangent kernel at initialization as
$K_{m}(X,X^{\prime})\;\coloneqq\;\big\langle\nabla_{\varphi}f\!\big(X;\varphi^{(0)}\big),\,\nabla_{\varphi}f\!\big(X^{\prime};\varphi^{(0)}\big)\big\rangle.$
By block structure, this decomposes as a sum over units and parameter blocks:
$K_{m}(X,X^{\prime})\;=\;\frac{1}{m}\sum_{i=1}^{m}\!\Biggl[\underbrace{\frac{\partial f}{\partial c_{i}}(X;\varphi^{(0)})\,\frac{\partial f}{\partial c_{i}}(X^{\prime};\varphi^{(0)})}_{\text{$c$–block}}\;+\;\underbrace{\big\langle\nabla_{U_{i}}f(X;\varphi^{(0)}),\,\nabla_{U_{i}}f(X^{\prime};\varphi^{(0)})\big\rangle}_{\text{$U$–block}}\\ \;+\;\underbrace{\big\langle\nabla_{W_{i}}f(X;\varphi^{(0)}),\,\nabla_{W_{i}}f(X^{\prime};\varphi^{(0)})\big\rangle_{F}}_{\text{$W$–block}}\Biggr].$
We remove the superscript and write $(c_{i}^{(0)},U_{i}^{(0)},W_{i}^{(0)})=(c_{i},U_{i},W_{i})$
cc–component.
$\frac{1}{m}\sum_{i=1}^{m}\frac{\partial f}{\partial c_{i}}(X;\varphi^{(0)})\,\frac{\partial f}{\partial c_{i}}(X^{\prime};\varphi^{(0)})$ $=\frac{1}{m}\sum_{i=1}^{m}\sigma\!\big({U_{i}}^{\top}a(X;W_{i})\big)\,\sigma\!\big(U_{i}^{\top}a(X^{\prime};W_{i})\big)$
$=\frac{2}{m}\sum_{i=1}^{m/2}\sigma\!\big({U_{i}}^{\top}a(X;W_{i})\big)\,\sigma\!\big(U_{i}^{\top}a(X^{\prime};W_{i})\big).$
Since the pairs $(c_{i},U_{i},W_{i})_{i=1}^{m/2}$ are i.i.d. at initialization, by the strong law of large numbers (SLLN),
$K_{c}(X,X^{\prime})\;:=\;\lim_{m\to\infty}\frac{2}{m}\sum_{i=1}^{m/2}\sigma\!\big(U_{i}^{\top}a(X;W_{i})\big)\,\sigma\!\big(U_{i}^{\top}a(X^{\prime};W_{i})\big)=\mathbb{E}\!\left[\,\sigma(U^{\top}a)\,\sigma(U^{\top}a^{\prime})\,\right].$
UU–component.
Writing $a_{i}=a(X;W_{i})$ and $a_{i}^{\prime}=a(X^{\prime};W_{i})$,
$\frac{1}{m}\sum_{i=1}^{m}\big\langle\nabla_{U_{i}}f(X;\varphi^{(0)}),\,\nabla_{U_{i}}f(X^{\prime};\varphi^{(0)})\big\rangle$ $=\frac{1}{m}\sum_{i=1}^{m}c_{i}^{2}\,\sigma^{\prime}\!\big(U_{i}^{\top}a_{i}\big)\,\sigma^{\prime}\!\big(U_{i}^{\top}a_{i}^{\prime}\big)\;\langle a_{i},a_{i}^{\prime}\rangle$
$=\frac{2}{m}\sum_{i=1}^{m/2}\sigma^{\prime}\!\big(U_{i}^{\top}a_{i}\big)\,\sigma^{\prime}\!\big(U_{i}^{\top}a_{i}^{\prime}\big)\;\langle a_{i},a_{i}^{\prime}\rangle$
where we used the fact that $c_{i}^{2}=1$ surely at initialization. Then SLLN yields
$K_{u}(X,X^{\prime})\;:=\;\lim_{m\to\infty}\frac{2}{m}\sum_{i=1}^{m/2}\sigma^{\prime}\!\big(U_{i}^{\top}a_{i}\big)\,\sigma^{\prime}\!\big(U_{i}^{\top}a_{i}^{\prime}\big)\;\langle a_{i},a_{i}^{\prime}\rangle=\mathbb{E}\!\left[\,\sigma^{\prime}(U^{\top}a)\,\sigma^{\prime}(U^{\top}a^{\prime})\,\langle a,a^{\prime}\rangle\,\right].$
WW–component.
Writing $M_{i}=X\,J_{s}(X^{\top}W_{i}q_{X})\,X^{\top}$ and $M_{i}^{\prime}=X^{\prime}\,J_{s}(X^{\prime\top}W_{i}q_{X^{\prime}})\,X^{\prime\top}$, using the Frobenius product identity $\langle xy^{\top},x^{\prime}y^{\prime\top}\rangle_{F}=\langle x,x^{\prime}\rangle\,\langle y,y^{\prime}\rangle$ and that $M$ and $M^{\prime}$ are symmetric,
$\frac{1}{m}\sum_{i=1}^{m}\big\langle\nabla_{W_{i}}f(X),\,\nabla_{W_{i}}f(X^{\prime})\big\rangle_{F}$ $=\frac{1}{m}\sum_{i=1}^{m}\sigma^{\prime}\!\big(U_{i}^{\top}a_{i}\big)\sigma^{\prime}\!\big(U_{i}^{\top}a_{i}^{\prime}\big)\big\langle(M_{i}U_{i})q_{X}^{\top},\;(M_{i}^{\prime}U_{i})q_{X^{\prime}}^{\top}\big\rangle_{F}$
$=\frac{2}{m}\sum_{i=1}^{m/2}\sigma^{\prime}\!\big(U_{i}^{\top}a_{i}\big)\sigma^{\prime}\!\big(U_{i}^{\top}a_{i}^{\prime}\big)\;\langle q_{X},q_{X^{\prime}}\rangle\;U_{i}^{\top}M_{i}M_{i}^{\prime}U_{i}.$
Again using SLLN,
$K_{w}(X,X^{\prime})\;:=\;\lim_{m\to\infty}\frac{2}{m}\sum_{i=1}^{m/2}\sigma^{\prime}\!\big(U_{i}^{\top}a_{i}\big)\sigma^{\prime}\!\big(U_{i}^{\top}a_{i}^{\prime}\big)\,\langle q_{X},q_{X^{\prime}}\rangle\,U_{i}^{\top}M_{i}M_{i}^{\prime}U_{i}=\langle q_{X},q_{X^{\prime}}\rangle\,\mathbb{E}\!\left[\sigma^{\prime}(U^{\top}a)\sigma^{\prime}(U^{\top}a^{\prime})\,U^{\top}MM^{\prime}U\right].$
Putting the three blocks together and taking $m\to\infty$,
$K(X,X^{\prime})\;=\;\lim_{m\to\infty}K_{m}(X,X^{\prime})\;=\;K_{c}(X,X^{\prime})+K_{u}(X,X^{\prime})+K_{w}(X,X^{\prime}).$
A.6 Proof of Lemma 5
Lemma 5.
Let $\tilde{f}(\cdot;v)\in\mathcal{F}_{\bar{\nu}}$ and $\tilde{\varphi}$ be as in (10). Define $\varepsilon_{\mathrm{app}}:=\sup_{j\in[n]}\ \big|f_{\mathrm{lin}}(X^{(j)};\tilde{\varphi})-\tilde{f}(X^{(j)};v)\big|$. Then for any $\delta\in(0,1)$, we have
$\varepsilon_{\mathrm{app}}\;\leq\;2(\sigma_{0}\bar{\nu}_{c}+\sigma_{1}\bar{\nu}_{u}+\sigma_{1}\bar{\nu}_{w})\sqrt{\tfrac{2\log(2n/\delta)}{m}}=:B_{\mathrm{app},m}(\delta),$
with probability at least $1-\delta$. We denote this event with $\mathcal{E}_{\mathrm{app}}(\delta)$.
Proof.
Fix $X\in\mathcal{X}$ and write $a(X;W)=:a$, $M(X;W)=:M$. The linear model evaluated at $\tilde{\varphi}$ decomposes as
$f_{\mathrm{lin}}(X;\tilde{\varphi})\;=\;\frac{1}{m}\sum_{i=1}^{m}g_{i}(X),$
where, with $\langle\cdot,\cdot\rangle_{F}$ the Frobenius product,
$g_{i}(X)$ $:=\phi_{c}\big(X;\varphi^{(0)}_{i}\big)\,v_{c}\!\big(\varphi^{(0)}_{i}\big)+\big\langle\phi_{u}\big(X;\varphi^{(0)}_{i}\big),\,v_{u}\!\big(\varphi^{(0)}_{i}\big)\big\rangle+\big\langle\phi_{w}\big(X;\varphi^{(0)}_{i}\big),\,v_{w}\!\big(\varphi^{(0)}_{i}\big)\big\rangle_{F},$ (19)
$\phi_{c}(X;\varphi_{0})$ $:=\sigma\!\big(U^{\top}a(X;W)\big),$
$\phi_{u}(X;\varphi_{0})$ $:=\sigma^{\prime}\!\big(U^{\top}a(X;W)\big)\,a(X;W),$
$\phi_{w}(X;\varphi_{0})$ $:=\sigma^{\prime}\!\big(U^{\top}a(X;W)\big)\,(M(X;W)\,U)\,q_{X}^{\top}.$
By the definition of $\tilde{f}(\cdot;v)$,
$\tilde{f}(X;v)\;:=\;\mathbb{E}\!\big[\phi_{c}(X;\varphi_{0})\,v_{c}(\varphi_{0})\big]+\mathbb{E}\!\big[\langle\phi_{u}(X;\varphi_{0}),v_{u}(\varphi_{0})\rangle\big]+\mathbb{E}\!\big[\langle\phi_{w}(X;\varphi_{0}),v_{w}(\varphi_{0})\rangle_{F}\big]\;=\;\mathbb{E}\!\big[g_{1}(X)\big],$ (20)
where the expectation is over a single draw $\varphi_{0}\sim\varphi^{(0)}_{1}$. Hence
$f_{\mathrm{lin}}(X;\tilde{\varphi})-\tilde{f}(X;v)\;=\;\frac{1}{m}\sum_{i=1}^{m}\Big(g_{i}(X)-\mathbb{E}[g_{i}(X)]\Big).$ (21)
Recall $m$ is even and $(\varphi^{(0)}_{i})_{i\leq m/2}$ are i.i.d., while $\varphi^{(0)}_{i+m/2}$ equals $\varphi^{(0)}_{i}$ with the $c_{i}$ sign-flipped. Define
$\bar{g}_{i}(X)\;:=\;\frac{g_{i}(X)+g_{i+m/2}(X)}{2},\qquad i=1,\dots,m/2.$
Then $\{\bar{g}_{i}(X)\}_{i=1}^{m/2}$ are i.i.d. and $\mathbb{E}\bar{g}_{1}(X)=\mathbb{E}g_{1}(X)$, so
$\Bigg|\frac{1}{m}\sum_{i=1}^{m}g_{i}(X)-\mathbb{E}[g_{1}(X)]\Bigg|\;=\;\Bigg|\,\frac{1}{m/2}\sum_{i=1}^{m/2}\bar{g}_{i}(X)-\mathbb{E}[\bar{g}_{1}(X)]\,\Bigg|.$ (22)
We use the Orlicz $\psi_{2}$ norm (see, e.g., (Vershynin, 2025, Chap. 2)):
$\|Z\|_{\psi_{2}}\;:=\;\inf\{t>0:\ \mathbb{E}\exp(Z^{2}/t^{2})\leq 2\}.$
(i) $c$ term. Since $|v_{c}(\varphi_{0})\phi_{c}|\leq\sigma_{0}\bar{\nu}_{c}$ a.s., then $\|\phi_{c}\,v_{c}\|_{\psi_{2}}\;\leq\;\sigma_{0}\,\bar{\nu}_{c}.$
(ii) $U$ term. Since $|\langle\phi_{u},v_{u}\rangle|\leq\|\phi_{u}\|_{2}\|v_{u}\|_{2}\leq\sigma_{1}\bar{\nu}_{u}$ a.s., we have $\|\langle\phi_{u},v_{u}\rangle\|_{\psi_{2}}\;\leq\;\sigma_{1}\,\bar{\nu}_{u}.$
(iii) $W$ term. We can write $\langle\phi_{w},v_{w}\rangle_{F}=\langle U,\sigma^{\prime}(U^{\top}a)M\,v_{w}(\varphi_{0})\,q_{X}\rangle$. Using the fact that $U$ is Gaussian and $\|\sigma^{\prime}(U^{\top}a)M\,v_{w}(\varphi_{0})\,q_{X}\|_{2}\leq\bar{\nu}_{w}$ a.s., we have $\|\langle\phi_{w},v_{w}\rangle_{F}\|_{\psi_{2}}\;\leq\;\sigma_{1}\;\sqrt{\tfrac{8}{3}}\;\bar{\nu}_{w}\;\leq\;2\sigma_{1}\bar{\nu}_{w}.$
Combining (i)–(iii) by the triangle inequality for $\psi_{2}$ norms,
$\|g_{i}(X)\|_{\psi_{2}}\;\leq\;K\quad\text{with}\quad K:=2(\sigma_{0}\bar{\nu}_{c}+\sigma_{1}\bar{\nu}_{u}+\sigma_{1}\bar{\nu}_{w}).$ (23)
Since $\{\bar{g}_{i}(X)\}_{i=1}^{m/2}$ are i.i.d. and centered after subtracting $\mathbb{E}\bar{g}_{1}(X)$, the standard sub-Gaussian tail bound yields, for all $t>0$,
$\mathbb{P}\!\left(\left|\frac{1}{m/2}\sum_{i=1}^{m/2}\big(\bar{g}_{i}(X)-\mathbb{E}\bar{g}_{1}(X)\big)\right|\geq t\right)\;\leq\;2\exp\!\left(-\,\frac{(m/2)\,t^{2}}{K^{2}}\right).$
Choosing $t\;=\;K\,\sqrt{\tfrac{2\log(2n/\delta)}{m}}$ and union bounding over $X^{(j)}$, we obtain with probability at least $1-\delta$,
$\sup_{j\in[n]}\Big|f_{\mathrm{lin}}(X^{(j)};\tilde{\varphi})-\tilde{f}(X^{(j)};v)\Big|\;\leq\;2\Big(\sigma_{0}\bar{\nu}_{c}+\sigma_{1}\bar{\nu}_{u}+\sigma_{1}\bar{\nu}_{w}\Big)\,\sqrt{\frac{2\log(2n/\delta)}{m}}.$
A.7 Proof of Theorem 1
The following lemma will be useful.
Lemma 6 (Uniform envelope bounds).
On the event $\mathcal{E}_{U,m}(\delta^{\prime})$ and for any $\varphi\in\Omega_{\rho}$,
$\sup_{X\in\mathcal{X}}|f(X;\varphi)|\;\leq\;L_{1,m}(\delta^{\prime})\,\sqrt{\rho_{u}^{2}+\rho_{w}^{2}}\;+\;\sigma_{0}\,\rho_{c}=:f^{\mathrm{max}}.$
Moreover, for any $\tilde{f}\in\mathcal{F}_{\bar{\nu}}$,
$\sup_{X\in\mathcal{X}}|\tilde{f}(X;v)|\;\leq\;\sigma_{0}\,\bar{\nu}_{c}\;+\;\sigma_{1}\,\bar{\nu}_{u}\;+\;\sigma_{1}\,\sqrt{d}\,\bar{\nu}_{w}=:y^{\mathrm{max}}.$
Proof.
Bounding $|f(X;\varphi)|$. Fix $X\in\mathcal{X}$ and write $\delta c_{i}=c_{i}-c_{i}^{(0)}$, $\delta\theta_{i}=(U_{i}-U_{i}^{(0)},\operatorname{vec}(W_{i}-W_{i}^{(0)}))$. Thanks to symmetric initialization, $f(X;\varphi^{(0)})=0$, hence
$|f(X;\varphi)|=\underbrace{\frac{1}{\sqrt{m}}\sum_{i=1}^{m}\Big|c_{i}^{(0)}\big(h(X;\theta_{i})-h(X;\theta_{i}^{(0)})\big)\Big|}_{(i)}+\underbrace{\frac{1}{\sqrt{m}}\sum_{i=1}^{m}\Big|\delta c_{i}\,h(X;\theta_{i})\Big|}_{(ii)}.$
On $\mathcal{E}_{U,m}(\delta^{\prime})$, Lemma 2 gives $\big|h(X;\theta_{i})-h(X;\theta_{i}^{(0)})\big|\leq L_{1,m}(\delta^{\prime})\|\delta\theta_{i}\|_{2}$ and we also have $|c_{i}^{(0)}|=1$ surely. Using Cauchy–Schwarz and $\varphi\in\Omega_{\rho}$, we get
$(i)\leq\frac{1}{\sqrt{m}}\sum_{i=1}^{m}L_{1,m}(\delta^{\prime})\|\delta\theta_{i}\|_{2}\;\leq\;L_{1,m}(\delta^{\prime})\sqrt{\rho_{u}^{2}+\rho_{w}^{2}}.$
Similarly, since we have $|h(X;\theta_{i})|\leq\sigma_{0}$, we get
$(ii)\;\leq\;\frac{1}{\sqrt{m}}\sum_{i=1}^{m}\frac{\rho_{c}}{\sqrt{m}}\sigma_{0}\;=\;\sigma_{0}\,\rho_{c}.$
Therefore, $|f(X;\varphi)|\leq L_{1,m}(\delta^{\prime})\sqrt{\rho_{u}^{2}+\rho_{w}^{2}}+\sigma_{0}\rho_{c}$, and taking the supremum over $X\in\mathcal{X}$ gives the stated bound.
Bounding $|\tilde{f}(X;v)|$. Recall
$\tilde{f}(X;v)=\mathbb{E}\!\left[\phi_{c}(X)\,v_{c}(\varphi_{0})\right]+\mathbb{E}\!\left[\langle\phi_{u}(X),v_{u}(\varphi_{0})\rangle\right]+\mathbb{E}\!\left[\langle\phi_{w}(X),v_{w}(\varphi_{0})\rangle_{F}\right],$
with $\phi_{c}(X)=\sigma(U^{\top}a)$, $\phi_{u}(X)=\sigma^{\prime}(U^{\top}a)\,a$, $\phi_{w}(X)=\sigma^{\prime}(U^{\top}a)\,(MU)\,q_{X}^{\top}$. Using $|\sigma|\leq\sigma_{0}$, $|\sigma^{\prime}|\leq\sigma_{1}$, $\|a\|_{2}\leq 1$, and $\|M\|_{F}\leq 1$, we obtain
$\big|\mathbb{E}[\phi_{c}(X)\,v_{c}(\varphi_{0})]\big|$ $\leq\sigma_{0}\,\bar{\nu}_{c},$
$\big|\mathbb{E}[\langle\phi_{u}(X),v_{u}(\varphi_{0})\rangle]\big|$ $\leq\sigma_{1}\,\bar{\nu}_{u},$
$\big|\mathbb{E}[\langle\phi_{w}(X),v_{w}(\varphi_{0})\rangle_{F}]\big|$ $\leq\sigma_{1}\,\mathbb{E}\!\left[\|(MU)q_{X}^{\top}\|_{F}\right]\bar{\nu}_{w}\;\leq\;\sigma_{1}\,\mathbb{E}\!\left[\|U\|_{2}\right]\bar{\nu}_{w}\;\leq\;\sigma_{1}\,\sqrt{d}\,\bar{\nu}_{w},$
where the last inequality uses $\mathbb{E}\|U\|_{2}\leq\sqrt{d}$ for $U\sim\mathcal{N}(0,I_{d})$. Summing the three bounds and taking the supremum over $X\in\mathcal{X}$ proves the claim.
Theorem 1.
Assume $f^{\star}\in\mathcal{F}_{\bar{\nu}}$ and the projection radius satisfies $\rho\geq\bar{\nu}$. Run ProjGD onto $\Omega_{\rho}$ for $\tau$ steps with the step size $\eta=1/\sqrt{\tau}$. Then, conditioned on the event $\mathcal{E}_{U,m}(\delta^{\prime})\cap\mathcal{E}_{\mathrm{app}}(\delta)$, we have
$\min_{s<\tau}\ \widehat{\mathcal{L}}_{n}\!\big(\varphi^{(s)}\big)\;\lesssim_{\bar{\sigma},\bar{\nu},\rho}\!\;\frac{L_{1,m}(\delta^{\prime})^{4}}{\sqrt{\tau}}\\ +L_{1,m}(\delta^{\prime})(B_{\mathrm{app},m}(\delta)+B_{\mathrm{lin},m}(\delta^{\prime})+\varepsilon_{\mathrm{CoF}}),$
and for the average iterate $\bar{\varphi}^{(\tau)}:=\frac{1}{\tau}\sum_{s=0}^{\tau-1}\varphi^{(s)}$,
$\widehat{\mathcal{L}}_{n}\!\big(\bar{\varphi}^{(\tau)}\big)\;\lesssim_{\bar{\sigma},\bar{\nu},\rho}\!\;\frac{L_{1,m}(\delta^{\prime})^{4}}{\sqrt{\tau}}\\ +L_{1,m}(\delta^{\prime})(B_{\mathrm{app},m}(\delta)+B_{\mathrm{lin},m}(\delta^{\prime})+\varepsilon_{\mathrm{CoF}})+B_{\mathrm{lin},m}(\delta^{\prime})^{2},$
where $B_{\mathrm{lin},m}(\delta^{\prime})$ and $B_{\mathrm{app},m}(\delta)$ are defined as in Lemmas 3 and 5, and
$\varepsilon_{\mathrm{CoF}}\;\lesssim_{\bar{\sigma},\bar{\nu},\rho}\;\frac{L_{1,m}(\delta^{\prime})+L_{2,m}(\delta^{\prime})}{\sqrt{m}}.$
Proof.
Set the Lyapunov function $\mathcal{V}(\varphi)=\|\varphi-\tilde{\varphi}\|_{2}^{2}$ and write the pre-projection point $\varphi^{\prime(s)}=\varphi^{(s)}-\eta\,\nabla\widehat{\mathcal{L}}_{n}(\varphi^{(s)})$. By non-expansivity of Euclidean projection onto a convex set and the Pythagorean identity,
$\mathcal{V}(\varphi^{(s+1)})\;\leq\;\mathcal{V}(\varphi^{\prime(s)})\;=\;\mathcal{V}(\varphi^{(s)})+2\eta\,\underbrace{\big\langle\nabla\widehat{\mathcal{L}}_{n}(\varphi^{(s)}),\,\tilde{\varphi}-\varphi^{(s)}\big\rangle}_{\text{(i)}}+\eta^{2}\,\underbrace{\big\|\nabla\widehat{\mathcal{L}}_{n}(\varphi^{(s)})\big\|_{2}^{2}}_{\text{(ii)}}.$ (24)
Controlling (i).
For shorthand write $f^{(s)}_{j}:=f(X^{(j)};\varphi^{(s)})$, $f^{\star}_{j}:=f^{\star}(X^{(j)})$ and $g^{(s)}_{j}:=\nabla_{\varphi}f(X^{(j)};\varphi^{(s)})$, $g^{(0)}_{j}:=\nabla_{\varphi}f(X^{(j)};\varphi^{(0)})$. Then
$\big\langle\nabla\widehat{\mathcal{L}}_{n}(\varphi^{(s)}),\,\tilde{\varphi}-\varphi^{(s)}\big\rangle$ $=\frac{2}{n}\sum_{j=1}^{n}\bigl(f^{(s)}_{j}-f^{\star}_{j}\bigr)\;\big\langle g^{(s)}_{j},\,\tilde{\varphi}-\varphi^{(s)}\big\rangle.$ (25)
Introduce the change-of-feature error
$\varepsilon^{\mathrm{CoF}}_{j}(\varphi^{(s)}):=\big\langle g^{(s)}_{j}-g^{(0)}_{j},\,\tilde{\varphi}-\varphi^{(s)}\big\rangle,$ (26)
and note that
$\big\langle g^{(s)}_{j},\,\tilde{\varphi}-\varphi^{(s)}\big\rangle$ $=\big\langle g^{(0)}_{j},\,\tilde{\varphi}-\varphi^{(s)}\big\rangle+\varepsilon^{\mathrm{CoF}}_{j}(\varphi^{(s)})$
$=f_{\mathrm{lin}}(X^{(j)};\tilde{\varphi})-f_{\mathrm{lin}}(X^{(j)};\varphi^{(s)})+\varepsilon^{\mathrm{CoF}}_{j}(\varphi^{(s)}).$ (27)
Now add and subtract $f^{\star}_{j}-f^{(s)}_{j}$ in (27) to obtain
$\big\langle g^{(s)}_{j},\,\tilde{\varphi}-\varphi^{(s)}\big\rangle=\underbrace{f_{\mathrm{lin}}(X^{(j)};\tilde{\varphi})-f^{\star}_{j}}_{\varepsilon^{\mathrm{app}}_{j}}+\bigl(f^{\star}_{j}-f^{(s)}_{j}\bigr)+\underbrace{f^{(s)}_{j}-f_{\mathrm{lin}}(X^{(j)};\varphi^{(s)})}_{\varepsilon^{\mathrm{lin}}_{j}(\varphi^{(s)})}+\varepsilon^{\mathrm{CoF}}_{j}(\varphi^{(s)}).$
From Lemmas 3 and 5 we have,
$\big|\varepsilon^{\mathrm{lin}}_{j}(\varphi^{(s)})\big|$ $\leq B_{\mathrm{lin},m}(\delta^{\prime}),$ (28)
$\big|\varepsilon^{\mathrm{app}}_{j}\big|$ $\leq B_{\mathrm{app},m}(\delta),$ (29)
To bound $\big|\varepsilon^{\mathrm{CoF}}_{j}(\varphi^{(s)})\big|$, we start from (26) and use the triangle and Cauchy–Schwarz inequalities,
[…]
$\leq\;\sum_{i=1}^{m}\big\|\nabla_{\varphi_{i}}f(X^{(j)};\varphi^{(s)})-\nabla_{\varphi_{i}}f(X^{(j)};\varphi^{(0)})\big\|_{2}\,\big\|\tilde{\varphi}_{i}-\varphi^{(s)}_{i}\big\|_{2}.$ (31)
By construction of $\tilde{\varphi}$ and $\Omega_{\rho}$,
$\big\|\tilde{\varphi}_{i}-\varphi^{(s)}_{i}\big\|_{2}\;\leq\;\frac{\|\bar{\nu}\|_{2}+\|\rho\|_{2}}{\sqrt{m}}.$ (32)
Set $\delta c_{i}:=c_{i}^{(s)}-c_{i}^{(0)}$, $\delta\theta_{i}:=\theta_{i}^{(s)}-\theta_{i}^{(0)}$. From Lemma 1,
$\nabla_{\theta_{i}}f(X;\varphi)\;=\;m^{-1/2}\,c_{i}\,\nabla_{\theta_{i}}h(X;\theta_{i}),\qquad\nabla_{c_{i}}f(X;\varphi)\;=\;m^{-1/2}\,h(X;\theta_{i}).$
Hence
$\nabla_{\theta_{i}}f\big(X^{(j)};\varphi^{(s)}\big)-\nabla_{\theta_{i}}f\big(X^{(j)};\varphi^{(0)}\big)$ $=\frac{c_{i}^{(0)}}{\sqrt{m}}\Big(\nabla_{\theta_{i}}h\big(X^{(j)};\theta_{i}^{(s)}\big)-\nabla_{\theta_{i}}h\big(X^{(j)};\theta_{i}^{(0)}\big)\Big)+\frac{\delta c_{i}}{\sqrt{m}}\,\nabla_{\theta_{i}}h\big(X^{(j)};\theta_{i}^{(s)}\big),$ (33)
$\nabla_{c_{i}}f\big(X^{(j)};\varphi^{(s)}\big)-\nabla_{c_{i}}f\big(X^{(j)};\varphi^{(0)}\big)$ $=\frac{1}{\sqrt{m}}\Big(h\big(X^{(j)};\theta_{i}^{(s)}\big)-h\big(X^{(j)};\theta_{i}^{(0)}\big)\Big).$ (34)
Invoking Lemma 2, we have for all $i$:
$\big\|\nabla_{\theta_{i}}h(X;\theta_{i}^{(s)})-\nabla_{\theta_{i}}h(X;\theta_{i}^{(0)})\big\|_{2}\leq L_{2,m}(\delta^{\prime})\,\|\delta\theta_{i}\|_{2},\qquad\big|h(X;\theta_{i}^{(s)})-h(X;\theta_{i}^{(0)})\big|\leq L_{1,m}(\delta^{\prime})\,\|\delta\theta_{i}\|_{2},$
and $\|\nabla_{\theta_{i}}h(X;\theta_{i}^{(s)})\|_{2}\leq L_{1,m}(\delta^{\prime})$. With $|c_{i}^{(0)}|=1$ at initialization, $|\delta c_{i}|\leq\rho_{c}/\sqrt{m}$, and $\|\delta\theta_{i}\|_{2}\leq\sqrt{\rho_{u}^{2}+\rho_{w}^{2}}/\sqrt{m}$, the bounds (33)–(34) yield
$\big\|\nabla_{\theta_{i}}f(X^{(j)};\varphi^{(s)})-\nabla_{\theta_{i}}f(X^{(j)};\varphi^{(0)})\big\|_{2}$ $\leq\frac{L_{2,m}(\delta^{\prime})}{\sqrt{m}}\,\frac{\sqrt{\rho_{u}^{2}+\rho_{w}^{2}}}{\sqrt{m}}\;+\;\frac{\rho_{c}}{\sqrt{m}}\,\frac{L_{1,m}(\delta^{\prime})}{\sqrt{m}}$
$=\frac{L_{2,m}(\delta^{\prime})\sqrt{\rho_{u}^{2}+\rho_{w}^{2}}+\rho_{c}L_{1,m}(\delta^{\prime})}{m},$ (35)
$\big\|\nabla_{c_{i}}f(X^{(j)};\varphi^{(s)})-\nabla_{c_{i}}f(X^{(j)};\varphi^{(0)})\big\|_{2}$ $\leq\frac{L_{1,m}(\delta^{\prime})}{\sqrt{m}}\,\frac{\sqrt{\rho_{u}^{2}+\rho_{w}^{2}}}{\sqrt{m}}=\frac{L_{1,m}(\delta^{\prime})\sqrt{\rho_{u}^{2}+\rho_{w}^{2}}}{m}.$ (36)
Combining the $c$- and $\theta$-blocks via $\|(x,y)\|_{2}\leq\|x\|_{2}+\|y\|_{2}$,
$\big\|\nabla_{\varphi_{i}}f(X^{(j)};\varphi^{(s)})-\nabla_{\varphi_{i}}f(X^{(j)};\varphi^{(0)})\big\|_{2}\;\leq\;\frac{(L_{1,m}(\delta^{\prime})+L_{2,m}(\delta^{\prime}))\sqrt{\rho_{u}^{2}+\rho_{w}^{2}}+\rho_{c}L_{1,m}(\delta^{\prime})}{m}.$ (37)
Plugging (32) and (37) into (30), we obtain, for every $j$,
$\big|\varepsilon^{\mathrm{CoF}}_{j}(\varphi^{(s)})\big|\;\leq\;\Big(\,(L_{1,m}(\delta^{\prime})+L_{2,m}(\delta^{\prime}))\sqrt{\rho_{u}^{2}+\rho_{w}^{2}}+\rho_{c}L_{1,m}(\delta^{\prime})\,\Big)\,\frac{\|\bar{\nu}\|_{2}+\|\rho\|_{2}}{\sqrt{m}}\;=:\;B_{\mathrm{CoF},m}(\delta^{\prime}).$ (38)
Putting (27) in (25) and using the bounds (28), (29) and (38), we have, for all $s\geq 0$,
$\big\langle\nabla\widehat{\mathcal{L}}_{n}(\varphi^{(s)}),\,\tilde{\varphi}-\varphi^{(s)}\big\rangle$ = $=-\,2\,\widehat{\mathcal{L}}_{n}(\varphi^{(s)})+\frac{2}{n}\sum_{j=1}^{n}\bigl(f^{(s)}_{j}-f^{\star}_{j}\bigr)\Big(\varepsilon^{\mathrm{app}}_{j}+\varepsilon^{\mathrm{lin}}_{j}(\varphi^{(s)})+\varepsilon^{\mathrm{CoF}}_{j}(\varphi^{(s)})\Big)$
 $\leq-\,2\,\widehat{\mathcal{L}}_{n}(\varphi^{(s)})+2\,(f^{\max}+y^{\max})\Big(B_{\mathrm{app},m}(\delta)+B_{\mathrm{lin},m}(\delta^{\prime})+B_{\mathrm{CoF},m}(\delta^{\prime})\Big),$ (39)
where $f^{\max}$ and $y^{\max}$ are defined as in Lemma 6.
Controlling (ii).
For any $j\in[n]$ and head $i\in[m]$,
$\Big\|\nabla_{\varphi_{i}}f\big(X^{(j)};\varphi^{(s)}\big)\Big\|_{2}\;\leq\;\frac{|h(X^{(j)};\theta_{i}^{(s)})|}{\sqrt{m}}+\frac{|c_{i}^{(s)}|}{\sqrt{m}}\,\Big\|\nabla_{\theta_{i}}h\big(X^{(j)};\theta_{i}^{(s)}\big)\Big\|_{2}\;\leq\;\frac{1}{\sqrt{m}}\Big(\sigma_{0}+|c_{i}^{(s)}|\,L_{1,m}(\delta^{\prime})\Big),$
and since $|c_{i}^{(s)}|\leq 1+\rho_{c}/\sqrt{m}$ on $\Omega_{\rho}$,
$\Big\|\nabla_{\varphi_{i}}f\big(X^{(j)};\varphi^{(s)}\big)\Big\|_{2}\;\leq\;\frac{1}{\sqrt{m}}\Big(\sigma_{0}+(1+\tfrac{\rho_{c}}{\sqrt{m}})\,L_{1,m}(\delta^{\prime})\Big).$ (40)
Summing the squares over $i=1,\dots,m$ gives
$\Big\|\nabla_{\varphi}f\big(X^{(j)};\varphi^{(s)}\big)\Big\|_{2}\;\leq\;A_{m}(\delta^{\prime})\qquad\text{with}\qquad A_{m}(\delta^{\prime}):=\sigma_{0}+\Big(1+\tfrac{\rho_{c}}{\sqrt{m}}\Big)L_{1,m}(\delta^{\prime}).$ (41)
Therefore,
$\Big\|\nabla\widehat{\mathcal{L}}_{n}(\varphi^{(s)})\Big\|_{2}$ $=\left\|\frac{2}{n}\sum_{j=1}^{n}\bigl(f^{(s)}_{j}-f^{\star}_{j}\bigr)\,\nabla_{\varphi}f\big(X^{(j)};\varphi^{(s)}\big)\right\|_{2}$
$\leq\frac{2}{n}\sum_{j=1}^{n}|f^{(s)}_{j}-f^{\star}_{j}|\,\Big\|\nabla_{\varphi}f\big(X^{(j)};\varphi^{(s)}\big)\Big\|_{2}\;\leq\;2\,(f^{\max}+y^{\max})\,A_{m}(\delta^{\prime}),$
and hence
$\Big\|\nabla\widehat{\mathcal{L}}_{n}(\varphi^{(s)})\Big\|_{2}^{2}\;\leq\;4\,(f^{\max}+y^{\max})^{2}\,A_{m}(\delta^{\prime})^{2}\;=:\;B_{\mathrm{grad},m}(\delta^{\prime}).$ (42)
Lyapunov drift.
Plugging (39) and (42) into (24) yields
$\mathcal{V}(\varphi^{(s+1)})-\mathcal{V}(\varphi^{(s)})\;\leq\;-4\eta\,\widehat{\mathcal{L}}_{n}(\varphi^{(s)})+4\eta\,(f^{\max}+y^{\max})\big(B_{\mathrm{app},m}(\delta)+B_{\mathrm{lin},m}(\delta^{\prime})+B_{\mathrm{CoF},m}(\delta^{\prime})\big)+\eta^{2}B_{\mathrm{grad},m}(\delta^{\prime}).$ (43)
Summing (43) over $s=0,\dots,\tau-1$ and using $\sum_{s<\tau}(\mathcal{V}(\varphi^{(s+1)})-\mathcal{V}(\varphi^{(s)}))=\mathcal{V}(\varphi^{(\tau)})-\mathcal{V}(\varphi^{(0)})$, we obtain after rearranging the terms
$\frac{1}{\tau}\sum_{s=0}^{\tau-1}\widehat{\mathcal{L}}_{n}(\varphi^{(s)})\;\leq\;\frac{\mathcal{V}(\varphi^{(0)})}{4\eta\tau}+(f^{\max}+y^{\max})\big(B_{\mathrm{app},m}(\delta)+B_{\mathrm{lin},m}(\delta^{\prime})+B_{\mathrm{CoF},m}(\delta^{\prime})\big)+\frac{\eta}{4}\,B_{\mathrm{grad},m}(\delta^{\prime}).$ (44)
Choosing $\eta=\tau^{-1/2}$ and using $\mathcal{V}(\varphi^{(0)})\leq\|\bar{\nu}\|_{2}^{2}$ gives
$\min_{0\leq s<\tau}\widehat{\mathcal{L}}_{n}\!\big(\varphi^{(s)}\big)\;\leq\;\frac{\|\bar{\nu}\|_{2}^{2}+B_{\mathrm{grad},m}(\delta^{\prime})}{4\sqrt{\tau}}+(f^{\max}+y^{\max})\big(B_{\mathrm{app},m}(\delta)+B_{\mathrm{lin},m}(\delta^{\prime})+B_{\mathrm{CoF},m}(\delta^{\prime})\big).$ (45)
Average iterate.
Define $\widehat{\mathcal{L}}_{n}^{\mathrm{lin}}(\varphi):=\frac{1}{n}\sum_{j=1}^{n}(f_{\mathrm{lin}}(X^{(j)};\varphi)-y^{(j)})^{2}$. From the uniform linearization bound in Lemma 3, for all $\varphi\in\Omega_{\rho}$,
$\widehat{\mathcal{L}}_{n}(\varphi)\leq 2\,\widehat{\mathcal{L}}_{n}^{\mathrm{lin}}(\varphi)+2\,B_{\mathrm{lin},m}(\delta^{\prime})^{2},\qquad\widehat{\mathcal{L}}_{n}^{\mathrm{lin}}(\varphi)\leq 2\,\widehat{\mathcal{L}}_{n}(\varphi)+2\,B_{\mathrm{lin},m}(\delta^{\prime})^{2}.$
By convexity of $\widehat{\mathcal{L}}_{n}^{\mathrm{lin}}$ and Jensen,
$\widehat{\mathcal{L}}_{n}^{\mathrm{lin}}\!\big(\bar{\varphi}^{(\tau)}\big)\;\leq\;\frac{1}{\tau}\sum_{s=0}^{\tau-1}\widehat{\mathcal{L}}_{n}^{\mathrm{lin}}\!\big(\varphi^{(s)}\big)\;\leq\;\frac{2}{\tau}\sum_{s=0}^{\tau-1}\widehat{\mathcal{L}}_{n}\!\big(\varphi^{(s)}\big)+2\,B_{\mathrm{lin},m}(\delta^{\prime})^{2}.$
Combining the two inequalities yields the “sandwich” inequality
$\widehat{\mathcal{L}}_{n}\!\big(\bar{\varphi}^{(\tau)}\big)\;\leq\;\frac{4}{\tau}\sum_{s=0}^{\tau-1}\widehat{\mathcal{L}}_{n}\!\big(\varphi^{(s)}\big)+6\,B_{\mathrm{lin},m}(\delta^{\prime})^{2}.$ (46)
Substituting (44) into (46) gives the desired inequality. ∎
A.8 Proof of Proposition 1
Proposition 1.
Under the same conditions as Theorem 1, we have
$\mathbb{E}\left[\min_{s<\tau}\widehat{\mathcal{L}}_{n}(\varphi^{(s)})\,|\,\varphi^{(0)}\right]\;\lesssim_{\bar{\sigma},\bar{\nu},\rho}\!\;\frac{L_{1,m}(\delta^{\prime})^{4}}{\sqrt{\tau}}+L_{1,m}(\delta^{\prime})(B_{\mathrm{app},m}(\delta)+B_{\mathrm{lin},m}(\delta^{\prime})+\varepsilon_{\mathrm{CoF}}),$
and for the average iterate,
$\mathbb{E}\left[\widehat{\mathcal{L}}_{n}\!\big(\bar{\varphi}^{(\tau)}\big)\,|\,\varphi^{(0)}\right]\!\;\lesssim_{\bar{\sigma},\bar{\nu},\rho}\!\;\frac{L_{1,m}(\delta^{\prime})^{4}}{\sqrt{\tau}}\\ +L_{1,m}(\delta^{\prime})(B_{\mathrm{app},m}(\delta)+B_{\mathrm{lin},m}(\delta^{\prime})+\varepsilon_{\mathrm{CoF}})+B_{\mathrm{lin},m}(\delta^{\prime})^{2}.$
Proof.
Define the Lyapunov function $\mathcal{V}(\varphi)=\|\varphi-\tilde{\varphi}\|_{2}^{2}$ and the pre‑projection iterate $\varphi^{\prime(s)}=\varphi^{(s)}-\eta\,\nabla_{\varphi}\ell_{J_{s}}(\varphi^{(s)})$. By non‑expansivity of projection and the Pythagorean identity,
$\mathcal{V}(\varphi^{(s+1)})\leq\mathcal{V}(\varphi^{(s)})+2\eta\underbrace{\langle\nabla_{\varphi}\ell_{J_{s}}(\varphi^{(s)}),\,\tilde{\varphi}-\varphi^{(s)}\rangle}_{\text{(i)}}+\eta^{2}\underbrace{\|\nabla_{\varphi}\ell_{J_{s}}(\varphi^{(s)})\|_{2}^{2}}_{\text{(ii)}}.$ (47)
Controlling (i).
Let $E_{s}[\cdot]:=\mathbb{E}_{J_{s}}[\cdot\,|\,J_{<s},\varphi^{(0)}]$ denote expectation w.r.t. the fresh index $J_{s}$, keeping the past $J_{<s}$ and $\varphi^{(0)}$ fixed (so $\varphi^{(s)}$ is fixed). Since $E_{s}\big[\nabla_{\varphi}\ell_{J_{s}}(\varphi^{(s)})\big]=\nabla\widehat{\mathcal{L}}_{n}(\varphi^{(s)})$, the same “Controlling (i)” computation as in the proof of Theorem 1 gives
$E_{s}\big[\langle\nabla_{\varphi}\ell_{J_{s}}(\varphi^{(s)}),\,\tilde{\varphi}-\varphi^{(s)}\rangle\big]\leq-\,2\,\widehat{\mathcal{L}}_{n}(\varphi^{(s)})+2\,(f^{\max}+y^{\max})\Big(B_{\mathrm{app},m}(\delta)+B_{\mathrm{lin},m}(\delta^{\prime})+B_{\mathrm{CoF},m}(\delta^{\prime})\Big).$
Controlling (ii).
Similarly, using the same arguments leading to (42), we have $\|\nabla_{\varphi}\ell_{J_{s}}(\varphi^{(s)})\|_{2}^{2}\leq B_{\mathrm{grad},m}(\delta^{\prime})$ almost surely and therefore
$E_{s}\big[\|\nabla_{\varphi}\ell_{J_{s}}(\varphi^{(s)})\|_{2}^{2}\big]\leq B_{\mathrm{grad},m}(\delta^{\prime}).$
Lyapunov drift.
Taking $E_{s}[\cdot]$ in (47) and combining the previous two steps yields
$E_{s}\big[\mathcal{V}(\varphi^{(s+1)})-\mathcal{V}(\varphi^{(s)})\big]\leq-4\eta\,\widehat{\mathcal{L}}_{n}(\varphi^{(s)})+4\eta\,(f^{\max}+y^{\max})\Big(B_{\mathrm{app},m}(\delta)+B_{\mathrm{lin},m}(\delta^{\prime})+B_{\mathrm{CoF},m}(\delta^{\prime})\Big)+\eta^{2}B_{\mathrm{grad},m}(\delta^{\prime}).$
Taking expectation over $J_{0},\ldots,J_{\tau-1}$, summing $s=0,\dots,\tau-1$, using $\sum_{s<\tau}\big(\mathcal{V}(\varphi^{(s+1)})-\mathcal{V}(\varphi^{(s)})\big)=\mathcal{V}(\varphi^{(\tau)})-\mathcal{V}(\varphi^{(0)})$, and $\mathcal{V}(\varphi^{(0)})\leq\|\bar{\nu}\|_{2}^{2}$, we obtain after rearranging the terms
$\frac{1}{\tau}\sum_{s=0}^{\tau-1}\mathbb{E}\big[\widehat{\mathcal{L}}_{n}(\varphi^{(s)})|\varphi^{(0)}\big]\leq\frac{\|\bar{\nu}\|_{2}^{2}}{4\eta\tau}+(f^{\max}+y^{\max})\Big(B_{\mathrm{app},m}(\delta)+B_{\mathrm{lin},m}(\delta^{\prime})+B_{\mathrm{CoF},m}(\delta^{\prime})\Big)+\frac{\eta}{4}B_{\mathrm{grad},m}(\delta^{\prime}).$ (48)
With $\eta=\tau^{-1/2}$ and $\min_{0\leq s<\tau}\widehat{\mathcal{L}}_{n}(\varphi^{(s)})\leq\tfrac{1}{\tau}\sum_{s<\tau}\widehat{\mathcal{L}}_{n}(\varphi^{(s)})$, we get
$\mathbb{E}\left[\min_{0\leq s<\tau}\widehat{\mathcal{L}}_{n}\!\big(\varphi^{(s)}\big)|\varphi^{(0)}\right]\;\leq\;\frac{\|\bar{\nu}\|_{2}^{2}+B_{\mathrm{grad},m}(\delta^{\prime})}{4\sqrt{\tau}}+(f^{\max}+y^{\max})\Big(B_{\mathrm{app},m}(\delta)+B_{\mathrm{lin},m}(\delta^{\prime})+B_{\mathrm{CoF},m}(\delta^{\prime})\Big).$ (49)
Average iterate.
Taking conditional expectation given $\varphi^{(0)}$ on both sides of (46), we get
$\mathbb{E}\!\left[\widehat{\mathcal{L}}_{n}\!\big(\bar{\varphi}^{(\tau)}\big)\,\middle|\,\varphi^{(0)}\right]\;\leq\;\frac{4}{\tau}\sum_{s=0}^{\tau-1}\mathbb{E}\!\left[\widehat{\mathcal{L}}_{n}\!\big(\varphi^{(s)}\big)\,\middle|\,\varphi^{(0)}\right]+6\,B_{\mathrm{lin},m}(\delta^{\prime})^{2}.$ (50)
Substituting (48) into (50) gives the desired inequality. ∎
Appendix B EXPERIMENTAL DETAILS
Code and Reproducibility.
All code to reproduce our experiments is available at the following repository:
https://github.com/enesarda22/nonasymptotic-transformer
B.1 IndRNN Architecture
Our recurrent model is an Independent RNN (IndRNN) studied by Cayci and Eryilmaz (2024a) with one key difference: because the prediction target in our case is a scalar, we use only the final hidden state and apply a linear layer, rather than producing a sequence of outputs. Concretely, for an input sequence $X=[X_{1},\ldots,X_{T}]\in\mathbb{R}^{d\times T},$ width $m$, activation $\sigma$, hidden matrix $U\in\mathbb{R}^{m\times d}$, diagonal recurrent weights $w\in\mathbb{R}^{m}$, and linear layer weights $c\in\mathbb{R}^{m}$, we define
$h_{0}$ $=0\in\mathbb{R}^{m},$ (51)
$h_{t}$ $=\sigma\!\big(U\,X_{t}+w\odot h_{t-1}\big)\in\mathbb{R}^{m},\qquad t=1,\ldots,T,$ (52)
$f_{\mathrm{rnn}}(X;\varphi_{\mathrm{rnn}})$ $:=\frac{1}{\sqrt{m}}\,c^{\top}h_{T}\in\mathbb{R},$ (53)
where $\odot$ denotes elementwise multiplication and $\varphi_{\mathrm{rnn}}=(\varphi_{\mathrm{rnn},i})_{i=1}^{m}$ with $\varphi_{\mathrm{rnn},i}:=(U_{i},w_{i},c_{i})$.
Symmetric initialization.
Similar to Section 5.2, for IndRNN, we use a symmetric initialization: assume $m$ is even and draw the first half independently as
$U_{i}^{(0)}\sim\mathcal{N}(0,I_{d}),\qquad w_{i}^{(0)}\sim\mathrm{Rad}(\gamma),\quad c_{i}^{(0)}\sim\mathrm{Rad}(1),\qquad i=1,\ldots,\tfrac{m}{2}.$ (54)
The second half is set as
$U_{i+m/2}^{(0)}=U_{i}^{(0)},\quad w_{i+m/2}^{(0)}=w_{i}^{(0)},\quad c_{i+m/2}^{(0)}=-\,c_{i}^{(0)}.$
This pairing ensures that the initial predictor is identically zero for any input $X\in\mathcal{X}$.
Parameter set.
Similar to Definition 2, for IndRNN, we define the neighborhood for projection to be
$\Omega_{\mathrm{rnn},\rho}:=\Big\{(U,w,c):\ \|U-U^{(0)}\|_{2,\text{row}}\leq\tfrac{\rho_{u}}{\sqrt{m}},\ \|w-w^{(0)}\|_{\infty}\leq\tfrac{\rho_{w}}{\sqrt{m}},\ \|c-c^{(0)}\|_{\infty}\leq\tfrac{\rho_{c}}{\sqrt{m}}\Big\}.$
Training error for IndRNN.
For completeness, we state the training loss result for the IndRNN architecture with scalar outputs, which follows as a corollary of (Cayci and Eryilmaz, 2024a, Thm. 4.4).
Proposition 2.
Assume $\tilde{f}\in\mathcal{F}_{\mathrm{rnn},\bar{\nu}}$ (see (Cayci and Eryilmaz, 2024a, Eq. (3.5))), and that the targets are generated as $y^{(j)}=\tilde{f}(X^{(j)})_{T}$ for $j=1,\dots,n$, with projection radius $\rho\geq\bar{\nu}$. Run ProjGD on $f_{\mathrm{rnn}}$ with projection onto $\Omega_{\mathrm{rnn},\rho}$ for $\tau$ steps with step size $\eta=1/\sqrt{\tau}$. Then, for any $\delta\in(0,1)$, with probability at least $1-\delta$,
$\min_{s<\tau}\ \widehat{\mathcal{L}}_{n}\!\big(\varphi_{\mathrm{rnn}}^{(s)}\big)\;\lesssim_{\bar{\sigma},\bar{\nu},\rho,\mu_{T}}\;\frac{1}{\sqrt{\tau}}\;+\;\sqrt{\frac{\log(2n/\delta)}{m}},$
where $\mu_{T}=\mathcal{O}(1)$ if $\gamma+\tfrac{\rho_{w}}{\sqrt{m}}\leq\tfrac{1}{\sigma_{1}}$, and $\mu_{T}=\exp(\Omega(T))$ otherwise. Here $\gamma>0$ is the Rademacher scale in the symmetric initialization of the diagonal recurrent weights (see (54)).
Max Jacobian norms in training.
For Transformer and IndRNN we calculate
$\mathcal{G}_{\mathrm{rnn}}(\varphi_{\mathrm{rnn}}):=\frac{1}{n}\sum_{j=1}^{n}\nabla_{w}f_{\mathrm{rnn}}(X^{(j)};\varphi_{\mathrm{rnn}}),\qquad\mathcal{G}_{\mathrm{tf}}(\varphi):=\frac{1}{n}\sum_{j=1}^{n}\nabla_{W}f(X^{(j)};\varphi).$
In Figure 2, we plot $\max_{s<\tau}\|\mathcal{G}_{\mathrm{rnn}}(\varphi_{\mathrm{rnn}}^{(s)})\|_{2}$ and $\max_{s<\tau}\|\mathcal{G}_{\mathrm{tf}}(\varphi^{(s)})\|_{F}$.